### v1

# On extension modules of finite homological dimension

Abstract.
We explore the implications of the finiteness of homological dimensions for Ext modules, focusing on projective dimension, injective dimension, and their Gorenstein counterpart. In this direction, we establish several finiteness criteria for homological dimensions. Our results include the consequences of the finiteness of the Gorenstein (injective) dimension of the deficiency modules as well as a duality for certain Ext modules, all of finite (Gorenstein) projective dimension.
Key words and phrases:
Finite homological dimension, Ext modules, Gorenstein ring, Deficiency module
2020 Mathematics Subject Classification:
Primary: 13D05, 13D07, 13H10. The second author was supported by grant 2019/21181-0, São Paulo Research Foundation (FAPESP). The third author was supported by grants 2022/03372-5 and 2023/15733-5, São Paulo Research Foundation (FAPESP)
1. Introduction
After the astonishing ideas of Auslander, Buchsbaum, and Serre culminating in the characterization of the regular property of a Noetherian local ring by the finiteness of the projective dimension of its residual field, homological algebra has played a fundamental role in commutative ring theory. Indeed, this foundation enabled diverse characterizations of ring-theoretic properties via finiteness conditions on homological dimensions (e.g., projective, complete intersection, Gorenstein dimensions, and their injective versions), see for instance some classical references [1, 2, 6, 15, 29]. Establishing criteria for the finiteness of homological dimensions as well as their prescription – especially via Ext vanishing – is a very active topic in homological commutative algebra, containing a vast literature [1, 4, 6, 9, 12, 13, 18, 19, 21, 23, 27, 28, 26, 37, 36].
Recently, an approach to study the implications of vanishing of Ext is under the condition that certain Hom module has some finite homological dimension, as evidenced in [11, 12, 20, 28]. In addition, Ghosh-Puthenpurakal [18] and Zargar-Gheibi [37] studied the consequences of having Ext modules with finite injective dimension, deriving criteria for finiteness for the projective dimension of modules. Motivated by all, in this paper, we aim to study the consequences of having Ext modules with some finite homological dimension, focusing on the projective dimension, injective dimension, and their Gorenstein versions. Thus, the standard assumption in the present text does not revolve around the usual Ext vanishing hypothesis but rather the finiteness of its homological dimensions.
Zargar and Gheibi proved in [37] that for non-zero finitely generated modules $M$ and $N$ over a local ring $R$ such that $\operatorname{id}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $i$, the finiteness of the projective dimension of $M$ is equivalent to that of injective dimension of $N$. Then, what can one tell if one consider the Ext modules $\operatorname{Ext}_{R}^{i\geq 0}(M,N)$ with another finite homological dimension under the additional condition that $\operatorname{pd}_{R}(M)<\infty$ or $\operatorname{id}_{R}(N)<\infty$? Based on their approach employed in [37, Section 4], we obtain results concerning the finiteness of the homological dimension of Ext of complexes (see Theorem 3.3 and Theorem 3.6), obtaining consequently the following:
Theorem 1.1 (Corollary 3.4 & Corollary 3.7).
Let $R$ be a local ring of dimension $d$, and let $M$ and $N$ be non-zero finitely generated $R$-modules. Let $\mathcal{H}$ be any of the homological dimensions $\operatorname{pd},\operatorname{G-dim},\operatorname{Gid}$, and suppose that $\mathcal{H}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq d-\operatorname{depth}(M)$.
  1. (1)
    Suppose that $\operatorname{pd}_{R}(M)<\infty$.
    1. (1.a)
      If $\mathcal{H}\in\{\operatorname{pd},\operatorname{G-dim}\}$, then $\mathcal{H}(N)<\infty$.
    2. (1.b)
      If $R$ has a dualizing complex and $\mathcal{H}=\operatorname{Gid}$, then $\operatorname{Gid}_{R}(N)<\infty$.
  2. (2)
    Suppose that $\operatorname{id}_{R}(N)<\infty$.
    1. (2.a)
      If $\mathcal{H}=\operatorname{pd}$, then $\operatorname{id}_{R}(M)<\infty$.
    2. (2.b)
      If $R$ has a dualizing complex and $\mathcal{H}=\operatorname{G-dim}$, then $\operatorname{Gid}_{R}(M)<\infty$.
    3. (2.c)
      If $\mathcal{H}=\operatorname{Gid}$, then $\operatorname{G-dim}_{R}(M)<\infty$.
Theorem 1.1 not only gives criteria for detecting the finiteness of homological dimension of modules but also allows one to obtain Gorenstein criteria as verified in Corollary 3.5 and Corollary 3.8. In addition, it is worth noting that Theorem 1.1 says that the counterparts with Gorenstein dimensions of the results of Foxby in [14, Proposition 3.5(b)] are true. Furthermore, it is worth to note that item (2.a) allows to recover the celebrated result of Foxby given in [15, Corollary 4.4], and that item (2.c) gives a positive answer to the Gorenstein version of the question in [18, Question 4.5].
In our study, we also consider the deficiency modules, which were introduced by Schenzel [33]. They are an example of Ext modules, so what can one tell if all of the deficiency modules of a module have some finite homological dimension? Fiel and the first author [13] proved that the finiteness of the projective (resp. injective) dimension of all the deficiency modules of a finitely generated module $M$ over a local ring $R$ implies that $M$ has finite injective (resp. projective) dimension. We not only prove this result with a different approach, but also that its Gorenstein version is true.
Theorem 1.2 (Theorem 4.1).
Let $R$ be a local ring with a dualizing complex and let $M$ be a finitely generated $R$-module.
  1. (1)
    If $\operatorname{G-dim}_{R}(K^{i}(M))<\infty$ for all $i\geq 0$, then $\operatorname{Gid}_{R}(M)<\infty$.
  2. (2)
    If $\operatorname{Gid}_{R}(K^{i}(M))<\infty$ for all $i\geq 0$, then $\operatorname{G-dim}_{R}(M)<\infty$.
Over Cohen-Macaulay local rings admitting canonical module, we furnish a duality between certain Ext modules under the finiteness of some homological dimensions. Such a duality is also established for semidualizing modules (Theorem 5.5) and is applied in many directions throughout the paper.
Theorem 1.3 (Theorem 5.6).
Assume $R$ Cohen-Macaulay local with canonical module $\omega$, and let $n$ be a non-negative integer. Let $\operatorname{H-dim}_{R}$ denote either the projective dimension or the Gorenstein dimension, and $\operatorname{Hid}_{R}$ the injective counterpart. Let $M$ and $N$ be finitely generated $R$-modules such that $\operatorname{H-dim}_{R}(N)<\infty$. Then the following conditions are equivalent:
  1. (1)
    $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq n$.
  2. (2)
    $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega))<\infty$ for all $0\leq i\leq n$.
In any case, we have natural isomorphisms
$\operatorname{Ext}_{R}^{i}(M,N)\otimes_{R}\omega\cong\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega)$
for all $0\leq i\leq n$.
This result leads to new finiteness criteria of homological dimension, for instance, we prove that if $R$ is a Cohen-Macaulay local ring of depth $t$ and, $M$ and $N$ are non-zero finitely generated $R$-modules such that $\operatorname{pd}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t-\operatorname{depth}(M)$, then finiteness of the projective dimension $M$ is equivalent to the finiteness to that of $N$, see Corollary 5.9. Moreover, our theorem allows us to show that over Cohen-Macaulay local rings, the Ghosh-Puthenpurakal type formula [18, Theorem 3.5] is true in other classes of modules than G-perfect modules, and, consequently, furnishes a Gorensteinness criterion for such a ring, see Corollary 5.11.
The organization of this paper is as follows. In Section 2, since our study considers tools from the derived category, we give preliminaries about it. In Section 3, we establish results concerning Ext of complexes of finite homological dimension, prove Theorem 1.1, and derive Gorenstein criteria. In Section 4, we study the implications of having deficiency modules with some finite homological dimension. In Section 5, we focus our study on modules over Cohen-Macaulay local rings, prove Theorem 5.6, and derive several consequences. In Section 6, we discuss some questions motivated by results in Section 5.
Next, we establish our setup and provide simple examples of $\operatorname{Ext}$ modules with finite homological dimension to motivate our study.
Setup.
Throughout this paper, $R$ stands for a Noetherian (commutative unitary) ring. Whenever we assume that $R$ is local, its maximal ideal will be denoted by $\mathfrak{m}$ and its residue field by $k$; also, its depth is denoted by $t$. In that case, for a finitely generated $R$-module $M$, its $i$-th syzygy module and its transpose of Auslander are considered defined by a minimal free resolution of $M$, and denoted by $\Omega^{i}(M)$ and $\operatorname{Tr}_{R}(M)$, respectively. Also, for an $R$-module $M$, we let $\operatorname{H-dim}_{R}(M)$ stand for $\operatorname{pd}_{R}(M)$ or $\operatorname{G-dim}_{R}(M)$, and $\operatorname{Hid}_{R}(M)$ for the corresponding injective version. For any unexplained terminology, we refer to [5] and [10].
Example 1.4.
Assume $R$ is a local ring and let $\underline{x}=x_{1},\ldots,x_{n}$ be a regular sequence. Let $N$ be a finitely generated $R$-module such that $\underline{x}N=0$. Set $M=R/(\underline{x})$. Since the minimal free resolution of $M$ is the Koszul complex of the sequence $\underline{x}$, we have
$\operatorname{Ext}^{i}_{R}(M,N)\cong N^{\binom{n}{i}},\quad\forall\ i=0,\ldots,n.$
Thus, $\operatorname{H-dim}_{R}(\operatorname{Ext}^{i}_{R}(M,N))<\infty$ for all $0\leq i\leq n$ if and only if $\operatorname{H-dim}_{R}(N)<\infty.$ Similarly, $\operatorname{Hid}_{R}(\operatorname{Ext}^{i}_{R}(M,N))<\infty$ for all $0\leq i\leq n$ if and only if $\operatorname{Hid}_{R}(N)<\infty.$
Now, set $L=\oplus_{i=1}^{n}R/(x_{1},\ldots,x_{i})$ and let $X$ be a finitely generated $R$-module such that $\underline{x}$ is $X$-regular. For all $i=1,\ldots,n$, by the Koszul duality we must have
$\operatorname{Ext}^{i}_{R}(L,X)\cong\operatorname{Ext}^{i}_{R}(R/(x_{1},\ldots,x_{i}),X)\cong X/(x_{1},\ldots,x_{i})X$
Thus $\operatorname{H-dim}_{R}(X)<\infty$ if and only if $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(L,X))<\infty$ for all $0\leq i\leq n$, and in that case, $\operatorname{H-dim}_{R}(\operatorname{Ext}^{i}_{R}(L,R))=i$, by the Auslander-Buchsbaum and Auslander-Bridge formulas.
Let $M$ be a finitely generated $R$-module. The grade of a finitely generated $R$-module $M$ is defined as $\operatorname{grade}_{R}(M)=\inf\{i\geq 0:\operatorname{Ext}_{R}^{i}(M,R)\not=0\}$. The module $M$ is said to be $\operatorname{H-}$perfect if $\operatorname{H-dim}_{R}(M)<\infty$ and $\operatorname{grade}(M)=\operatorname{H-dim}_{R}(M)$.
Example 1.5.
Let $M_{1},\ldots,M_{n}$ be H-perfect $R$-modules of grade $g_{1},\ldots,g_{n}$, respectively, and set $M=\bigoplus_{i=1}^{n}M_{i}$. Then
$\operatorname{Ext}^{g}_{R}(M,R)\cong\begin{cases}\operatorname{Ext}^{g_{i}}_{R}(M_{i},R),&\quad g=g_{i}\\ 0,&\quad\mbox{else}\end{cases}$
and $\operatorname{H-dim}_{R}(\operatorname{Ext}^{g_{i}}_{R}(M,R))=g_{i}$ by [3, Theorem 6.3].
An $R$-module $M$ is $p$-spherical if $\operatorname{pd}_{R}(M)\leq p$ and $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $11\leq i\leq p-$ 1.
Example 1.6.
Based on the proof of [25, Proposition 6] we can construct spherical non-perfect modules as follows. We repeat the proof for convenience. Let $M$ be a finitely generated $R$-module with positive grade $n$, and let $F_{\bullet}$ be a free resolution of $M$. Consider the truncation
[…]
Set $N:=\operatorname{coker}(\phi^{*})$. Since $\operatorname{grade}(M)=n>0$, the complex
[…]
is a free resolution of $N$ and
[…]
is exact. This means that $N$ is $n$-spherical as well as a $(n+1)$th syzygy of $M$ (in the sense of [26]), and also $\operatorname{Ext}^{n}_{R}(N,R)\cong M$. It follows that whenever $M$ is such that $\operatorname{pd}_{R}(M)>\operatorname{grade}(M)=n>0$, there exists a $n$-spherical non-perfect $R$-module $N$ such that $\operatorname{Ext}^{n}_{R}(N,R)\cong M$. In particular, if $\operatorname{H-dim}_{R}(M)<\infty$ (resp. $\operatorname{Hid}_{R}(M)<\infty)$ then $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{n}(N,R))<\infty$ (resp. $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{n}(N,R)<\infty$).
It is noteworthy that the previous procedure works likewise if one considers resolutions by totally reflexive modules instead. In this case, $N$ would still be a non-G-perfect module which is a $(n+1)$th syzygy of $M$ (again in the sense of [26]) such that $\operatorname{G-dim}_{R}(N)\leq n$ and $\operatorname{Ext}^{i}_{R}(N,R)=0$ for all $i\neq 0,n$.
Example 1.7.
Let $k$ be a field and consider $R=k[[x,y]]/(xy)$. Then $M=R/(x)$ has infinite free resolution
[…]
so that $\operatorname{Ext}^{2i+1}_{R}(M,M)\cong M$ for all $i$, which means that $\operatorname{CI-dim}_{R}(\operatorname{Ext}^{2i+1}_{R}(M,M))=0$ and $\operatorname{pd}_{R}(\operatorname{Ext}^{2i+1}_{R}(M,M))=\infty$.
2. Preliminaries on Derived Categories
In this section, we give some preliminaries concerning the derived category. For unexplained terminology, we refer to [6, Appendix], and for background on derived category theory, we additionally refer to [10].
2.1. Complexes and morphisms
All complexes considered in this section are descending. Let
$\boldsymbol{X}:\cdots\longrightarrow X_{i+1}\stackrel{{\scriptstyle\partial_{i+1}^{\boldsymbol{X}}}}{{\longrightarrow}}X_{i}\stackrel{{\scriptstyle\partial_{i}^{\boldsymbol{X}}}}{{\longrightarrow}}X_{i-1}\longrightarrow\cdots$
be a complex of $R$-modules.
If $X_{i}=0$ for all $i\gg 0$ (resp. for all $i\ll 0$), then $\boldsymbol{X}$ is said to be bounded to the left (resp. bounded to the right). If $\boldsymbol{X}$ is a bounded to the right and to the left, it is said to be bounded. We say that $\boldsymbol{X}$ is homologically bounded to the left (resp. homologically bounded to the right) if $\operatorname{H}_{i}(\boldsymbol{X})=0$ for all $i\gg 0$. If $\boldsymbol{X}$ is homologically bounded to the left and to the right, then $\boldsymbol{X}$ is said to be homologically bounded.
We recall that an morphism $f:\boldsymbol{X}\rightarrow\boldsymbol{Y}$ of complexes is a family $\{f_{i}:X_{i}\to Y_{i}\}_{i\in\mathbb{Z}}$ of homomorphism of $R$-modules such that $\partial_{i}^{\boldsymbol{Y}}\circ f_{i}=f_{i-1}\circ\partial_{i}^{\boldsymbol{X}}$. It is natural to define the composition of morphisms of complexes. The category of complexes over $R$ and homomorphism of $R$-modules is denoted by $\mathcal{C}(R)$.
Note that a morphism $f:\boldsymbol{X}\rightarrow\boldsymbol{Y}$ of complexes induces naturally homomorphisms $\operatorname{H}_{i}(f):\operatorname{H}_{i}(\boldsymbol{X})\rightarrow\operatorname{H}_{i}(\boldsymbol{Y})$. We say that $f$ is said to be a quasi-isomorphism if all the induced maps $H_{i}(f)$ are isomorphisms.
[…]
An $R$-module $M$ is identified naturally as a complex $\boldsymbol{X}$ such that $X_{0}=M$ and $X_{i}=0$ for all $i\neq 0$.
2.2. Notation of some full subcategories of $\mathcal{C}(R)$
We are considering the following full subcategories of $\mathcal{C}(R)$:
  1. (1)
    $\mathcal{C}_{\sqsubset}(R)$ (resp. $\mathcal{C}_{\sqsupset}(R)$) denotes the category of complexes bounded to the left (resp. right).
  2. (2)
    $\mathcal{C}_{(\sqsubset)}(R)$ (resp. $\mathcal{C}_{(\sqsupset)}(R)$) denotes the category of homologically bounded complexes to the left (resp. to the right).
  3. (3)
    $\mathcal{C}_{(\square)}(R)$ denotes the category of homologically bounded complexes.
  4. (4)
    $\mathcal{C}^{\rm P}(R)$ (resp. $\mathcal{C}^{\rm I}(R)$) denotes the category of complexes of $R$-modules of projective modules (resp. injective modules).
  5. (5)
    $\mathcal{C}^{(\rm f)}(R)$ denote the category of complexes with finitely generated homologies.
Moreover, for a subcategory $\mathcal{S}(R)\subseteq\mathcal{C}(R)$ we set $\mathcal{S}^{(\rm f)}(R)=\mathcal{S}(R)\cap\mathcal{C}^{(\rm f)}(R)$.
2.3. Shift
Let $n$ be an integer. For a complex $\boldsymbol{X},$ its $n$-fold shift $\Sigma^{n}\boldsymbol{X}$ is defined by:
$(\Sigma^{n}\boldsymbol{X})_{i}=X_{i-n}\,\,\text{ and }\,\,\partial_{i}^{\Sigma^{n}\boldsymbol{X}}=(-1)^{n}\partial_{i-n}^{\boldsymbol{X}}.$
2.4. Hom and tensor of complexes
Let $\boldsymbol{X}$ and $\boldsymbol{Y}$ be complexes of $R$-modules. The Hom complex $\operatorname{Hom}_{R}(\boldsymbol{X},\boldsymbol{Y})$ is defined by:
$\operatorname{Hom}_{R}(\boldsymbol{X},\boldsymbol{Y})_{i}=\prod_{j\in\mathbb{Z}}\operatorname{Hom}_{R}(X_{j},Y_{i+j})$
and
$\partial_{i}^{\text{Hom}_{R}(\boldsymbol{X},\boldsymbol{Y})}((\psi_{j}))=(\partial_{j}^{\boldsymbol{Y}}\circ\psi_{j}-(-1)^{i}\psi_{p-1}\circ\partial_{j}^{\boldsymbol{X}})_{j}$
The tensor product complex $\boldsymbol{X}\otimes_{R}\boldsymbol{Y}$ is defined by:
$\left(\boldsymbol{X}\otimes_{R}\boldsymbol{Y}\right)_{i}=\bigoplus_{j\in\mathbb{Z}}X_{j}\otimes_{R}Y_{i-j}$
and
$\partial_{i}^{\boldsymbol{X}\otimes_{R}\boldsymbol{Y}}\left(m_{j}\otimes n_{i-j}\right)=\partial_{i}^{\boldsymbol{X}}\left(m_{j}\right)\otimes n_{i-j}+(-1)^{j}m_{j}\otimes\partial_{i-j}^{\boldsymbol{Y}}\left(n_{i-j}\right).$
2.5. Derived category
We denote by $\mathcal{D}(R)$ the derived category of $R$. This is the category of $R$-complexes localized at the class of all quasi-isomorphisms. Describing this category can be technical, so for more details, we refer the reader to [10]. Nevertheless, we mention the following facts:
  1. (1)
    The objects of $\mathcal{D}(R)$ are exactly the objects of $\mathcal{C}(R)$.
  2. (2)
    A morphism in $\mathcal{C}(R)$ is a morphism in $\mathcal{D}(R)$, and it is an isomorphism in $\mathcal{D}(R)$ if and only if it is a quasi-isomorphism in $\mathcal{C}(R)$.
By $\simeq$ we indicate isomorphism in $\mathcal{D}(R)$.
2.6. Derived functors
The symbols $\operatorname{\mathbf{R}Hom}(-,-)$ and $-\otimes_{R}^{\mathbf{L}}-$ denote, respectively, the right and left derived functors of the homomorphism and tensor product functors for complexes over $R$. They are bifunctors on the derived category of $R$, defined up to isomorphism in $\mathcal{D}(R)$.
The bifunctor $\operatorname{\mathbf{R}Hom}(-,-)$ satisfies the following: Let $\boldsymbol{X}$ and $\boldsymbol{Y}$ be complexes.
  1. (1)
    If $\boldsymbol{X}\in\mathcal{C}_{(\sqsupset)}(R)$ and if $\boldsymbol{P}$ is a complex in $\mathcal{C}_{\sqsupset}^{\rm P}(R)$ such that $\boldsymbol{X}\simeq\boldsymbol{P}$, then $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{X},\boldsymbol{Y})=\operatorname{Hom}_{R}(\boldsymbol{P},\boldsymbol{Y})$.
  2. (2)
    If $\boldsymbol{Y}\in\mathcal{C}_{(\sqsubset)}(R)$ and if $\boldsymbol{I}$ is a complex in $\operatorname{C}_{\sqsubset}^{\rm I}(R)$ such that $\boldsymbol{I}\simeq\boldsymbol{Y}$, then $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{X},\boldsymbol{Y})=\operatorname{Hom}_{R}(\boldsymbol{X},\boldsymbol{I})$.
Further, the bifunctor $-\otimes_{R}^{\mathbf{L}}-$ satisfies the following: Let $\boldsymbol{X}$ and $\boldsymbol{Y}$ be complexes.
  1. (1)
    If $\boldsymbol{X}\in\mathcal{C}_{(\sqsupset)}(R)$ and if $\boldsymbol{P}$ is a complex in $\mathcal{C}_{\sqsupset}^{\rm P}(R)$ such that $\boldsymbol{X}\simeq\boldsymbol{P}$ , then $\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\boldsymbol{Y}=\boldsymbol{P}\otimes_{R}\boldsymbol{Y}$.
  2. (2)
    If $\boldsymbol{Y}\in\mathcal{C}_{(\sqsupset)}(R)$ and if $\boldsymbol{P}$ is a complex in $\mathcal{C}_{\sqsupset}^{\rm P}(R)$ such that $\boldsymbol{Y}\simeq\boldsymbol{P}$, then $\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\boldsymbol{Y}=\boldsymbol{X}\otimes_{R}\boldsymbol{P}$.
If $\boldsymbol{X}$ and $\boldsymbol{Y}$ are complexes, we set
$\operatorname{Ext}_{R}^{i}(\boldsymbol{X},\boldsymbol{Y})=\operatorname{H}_{-i}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{X},\boldsymbol{Y}))$ and $\operatorname{Tor}_{i}^{R}(\boldsymbol{X},\boldsymbol{Y})=\operatorname{H}_{i}(\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\boldsymbol{Y}),\text{ for all }i.$
2.7. Standard isomorphisms
Let $\boldsymbol{X},\boldsymbol{Y}$ and $\boldsymbol{Z}$ be complexes. Then we have the following isomorphisms in $\mathcal{D}(R)$:
  1. (Com)
    $\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\boldsymbol{Y}\simeq\boldsymbol{Y}\otimes_{R}^{\mathbf{L}}\boldsymbol{L}$.
  2. (Asoc)
    $(\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\boldsymbol{Y})\otimes_{R}^{\mathbf{L}}\boldsymbol{Z}\simeq\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\left(\boldsymbol{Y}\otimes_{R}^{\mathbf{L}}\boldsymbol{Z}\right).$
  3. (Adj)
    $\operatorname{\mathbf{R}Hom}_{R}\left(\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\boldsymbol{Y},\boldsymbol{Z}\right)\simeq\operatorname{\mathbf{R}Hom}_{R}\left(\boldsymbol{X},\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{Y},\boldsymbol{Z})\right).$
  4. (Swap)
    $\operatorname{\mathbf{R}Hom}_{R}\left(\boldsymbol{X},\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{Y},\boldsymbol{Z})\right)\simeq\operatorname{\mathbf{R}Hom}_{R}\left(\boldsymbol{Y},\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{X},\boldsymbol{Z})\right).$
2.8. Homological dimensions
(1) The projective dimension of a complex $\boldsymbol{X}\in\mathcal{C}_{(\sqsupset)}(R)$ is defined as:
$\operatorname{pd}_{R}(\boldsymbol{X})=\inf\{\sup\{i\in\mathbb{Z}:P_{i}\not=0\}:\boldsymbol{X}\simeq\boldsymbol{P}\in\mathcal{C}_{\sqsupset}^{\rm P}(R)\}.$
(2) The injective dimension of a complex $\boldsymbol{Y}\in\mathcal{C}_{(\sqsupset)}(R)$ is defined as:
$\operatorname{id}_{R}(\boldsymbol{Y})=\inf\{\sup\{i\in\mathbb{Z}:I_{-i}\not=0\}:\boldsymbol{Y}\simeq\boldsymbol{I}\in\mathcal{C}_{\sqsubset}^{\rm I}(R)\}.$
(3) A finitely generated $R$-module $G$ is totally reflexive if $G$ is reflexive and $\operatorname{Ext}_{R}^{i}(G,R)=\operatorname{Ext}_{R}^{i}(G^{\ast},R)=0$ for all $i>0$. By $\mathcal{C}^{\operatorname{G}}(R)$ we denote the category of complexes of totally reflexives $R$-modules. The Gorenstein dimension of a complex $\boldsymbol{X}\in\mathcal{C}_{(\sqsupset)}^{({\rm f})}(R)$ is defined as
$\operatorname{G-dim}_{R}(\boldsymbol{X})=\inf\{\sup\{i\in\mathbb{Z}:G_{i}\not=0\}:\boldsymbol{X}\simeq\boldsymbol{G}\in\mathcal{C}_{\sqsupset}^{\rm G}(R)\}.$
(4) An $R$-module $M$ is said to be $G$-injective if there exists an exact complex of injective $R$-modules such that
$\boldsymbol{I}:\cdots\xrightarrow{d_{2}}{I_{2}}\xrightarrow{d_{1}}{I_{0}}\xrightarrow{d_{0}}I_{-1}\xrightarrow{d-1}\cdots$
such that $M\cong\ker(d_{0})$ and the complex $\operatorname{Hom}_{R}(J,\boldsymbol{I})$ is also exact for any injective $R$-module $J$. By $\mathcal{C}^{\operatorname{GI}}(R)$ we denote the category of complexes of Gorenstein injective modules. The Gorenstein injective dimension of a complex $\boldsymbol{Y}\in\mathcal{C}_{(\sqsubset)}(R)$ is defined as:
$\operatorname{Gid}_{R}(\boldsymbol{Y})=\inf\{\sup\{i\in\mathbb{Z}:I_{-i}\not=0\}:\boldsymbol{Y}\simeq\boldsymbol{I}\in\mathcal{C}_{\sqsubset}^{\rm GI}(R)\}.$
Theorem 2.1.
The following class of $\mathcal{C}(R)$ are closed under short exact sequences.
  1. (1)
    ([10, Corollary 8.1.9]) The class of complexes in $\mathcal{C}_{(\sqsupset)}(R)$ with finite projective dimension.
  2. (2)
    ([10, Corollary 8.2.9])  The class of complexes in $\mathcal{C}_{(\sqsubset)}(R)$ with finite injective dimension.
  3. (3)
    ([6, Lemma 2.1.12 & Corollary 2.3.8]) The class of complexes in $\mathcal{C}_{(\sqsupset)}^{(\rm f)}(R)$ with finite Gorenstein dimension.
  4. (4)
    ([6, Lemma 3.2.12 & Corollary 6.2.5])  The class of complexes in $\mathcal{C}_{(\sqsubset)}(R)$ with finite Gorenstein injective dimension.
Theorem 2.2.
Let $R$ be a local ring and $\boldsymbol{X}$ be a complex.
  1. (1)
    ([10, Proposition 18.3.9]) If $\boldsymbol{X}\in\mathcal{C}_{(\sqsupset)}^{(\rm f)}(R)$, then $\operatorname{pd}_{R}(\boldsymbol{X})=\operatorname{pd}_{\widehat{R}}(\boldsymbol{X}\otimes_{R}\widehat{R}).$
  2. (2)
    ([10, Proposition 18.3.10]) If $\boldsymbol{X}\in\mathcal{C}_{(\sqsubset)}^{(\rm f)}(R)$, then $\operatorname{id}_{R}(\boldsymbol{X})=\operatorname{id}_{\widehat{R}}(\boldsymbol{X}\otimes_{R}\widehat{R}).$
  3. (3)
    ([23, 2.4]) If $\boldsymbol{X}\in\mathcal{C}_{(\square)}^{(\rm f)}(R)$, then $\operatorname{G-dim}_{R}(\boldsymbol{X})=\operatorname{G-dim}_{\widehat{R}}(\boldsymbol{X}\otimes_{R}\widehat{R}).$
  4. (4)
    ([10, Theorem 19.2.34]) If $\boldsymbol{X}\in\mathcal{C}_{(\square)}^{(\rm f)}(R)$ and $\operatorname{Gid}_{R}(\boldsymbol{X})<\infty$, then $\operatorname{Gid}_{R}(\boldsymbol{X})=\operatorname{Gid}_{\widehat{R}}(\boldsymbol{X}\otimes_{R}\widehat{R}).$
Proposition 2.3 ([4, Theorem 8.7]).
Let $R$ be a local ring and let $M$ be a finitely generated $R$-module. Then:
  1. (1)
    If $\operatorname{H\text{-}dim}_{R}(M)<\infty$, then $\operatorname{H\text{-}dim}_{R}(M)+0pt(M)=0pt(R)$.
  2. (2)
    For every integer $n\geq 0$, $\operatorname{H\text{-}dim}_{R}(\Omega^{n}(M))=\max\{\operatorname{H\text{-}dim}_{R}(M)-n,\,0\}.$
  3. (3)
    $\operatorname{G-dim}_{R}(M)\leq\operatorname{pd}_{R}(M)$ with equality whenever $\operatorname{pd}_{R}(M)$ is finite.
2.9. Semidualizing complexes.
An $R$-complex $\boldsymbol{C}\in\mathcal{C}_{(\square)}^{(\rm f)}(R)$ is semidualizing if the natural morphism $R\to\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{C},\boldsymbol{C})$ is an isomorphism in $\mathcal{D}(R)$. An $R$-module $C$ is said to be semidualizing if it corresponds to a semidualizing complex in $\mathcal{C}_{(\square)}^{(\rm f)}(R)$ ([7]).
2.10. Auslander categories
Let $\boldsymbol{C}$ be a semidualizing complex.
  1. (1)
    The Auslander class of $R$, denoted by $\mathcal{A}_{\boldsymbol{C}}(R)$, is the full subcategory of $\mathcal{D}(R)$ composed for all the complex $\boldsymbol{X}\in\mathcal{C}_{(\square)}(R)$ such that $\boldsymbol{C}\otimes_{R}^{\mathbf{L}}\boldsymbol{X}\in\mathcal{C}_{(\square)}(R)$ and the canonical map
    $\boldsymbol{X}\to\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{C},\boldsymbol{C}\otimes_{R}^{\mathbf{L}}\boldsymbol{X})$
    is an isomorphism in $\mathcal{D}(R)$.
  2. (2)
    The Bass class of $R$, denoted by $\mathcal{B}_{\boldsymbol{C}}(R)$, is the full subcategory of $\mathcal{C}_{(\square)}(R)$ composed for all the complexes $\boldsymbol{Y}\in\mathcal{C}_{(\square)}(R)$ such that $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{C},\boldsymbol{Y})\in\mathcal{C}_{(\square)}(R)$ and the canonical map
    $\boldsymbol{C}\otimes_{R}^{\mathbf{L}}\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{C},\boldsymbol{Y})\to\boldsymbol{Y}$
    is an isomorphism in $\mathcal{D}(R)$.
Denote by $\mathcal{F}(R)$ and $\mathcal{I}(R)$, respectively, the subcategories of $\mathcal{D}(R)$ consisting of complexes isomorphic to homologically bounded complexes of flat $R$-modules and of injective $R$-modules. Then $\mathcal{F}(R)\subset\mathcal{A}_{\boldsymbol{C}}(R)$ and $\mathcal{I}(R)\subset\mathcal{B}_{\boldsymbol{C}}(R)$.
2.11. Dualizing complex
Assume that $R$ is a local ring. A dualizing complex $\boldsymbol{D}$ is a bounded complex of injective $R$-modules with finitely generated homologies such that the natural morphism $R\rightarrow\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{D})$ is an isomorphism in $\mathcal{D}(R)$.
The ring $R$ has a dualizing complex if and only if $R$ is a homomorphic image of a Gorenstein local ring ([6, A.8.3.2] and [24, Corollary 1.4]). In particular, if $R$ is complete, then $R$ admits a dualizing complex. If $\boldsymbol{D}$ and $\boldsymbol{D}^{\prime}$ are dualizing complexes, then $\boldsymbol{D}\simeq\Sigma^{n}\boldsymbol{D}$ for some integer $n$ ([6, A.8.3.3]).
If $\boldsymbol{X}$ is a complex in $\mathcal{C}_{(\square)}^{(\rm f)}(R)$, then $\boldsymbol{X}\simeq\operatorname{\mathbf{R}Hom}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{X},\boldsymbol{D}),\boldsymbol{D})$ ([6, A.8.5]).
Proposition 2.4.
Assume $R$ is a local ring and let $\boldsymbol{D}$ be a dualizing complex.
  1. (1)
    ([6, Corollary 3.1.11]) A complex $\boldsymbol{X}$ in $\mathcal{C}_{(\sqsupset)}^{(\rm f)}(R)$ has finite Gorenstein dimension if and only if $\boldsymbol{X}\in\mathcal{A}_{\boldsymbol{D}}(R)$.
  2. (2)
    ([9, Theorem 4.4]) A complex $\boldsymbol{X}$ in $\mathcal{C}_{(\sqsubset)}(R)$ has finite Gorenstein injective dimension if and only if $\boldsymbol{X}\in\mathcal{B}_{\boldsymbol{D}}(R)$.
A celebrated result that relates to the Auslander categories, and the subcategories $\mathcal{F}(R)$ and $\mathcal{I}(R)$ is the following, which is known as the Foxby Equivalence.
Theorem 2.5 ([6, Theorem 3.3.2]).
Assume that $R$ is a local ring with a dualizing complex $\boldsymbol{D}$. Let $\boldsymbol{X}$ and $\boldsymbol{Y}$ be complexes in $\mathcal{C}_{(\square)}(R)$. Then:
$\boldsymbol{X}\in\mathcal{A}_{\boldsymbol{D}}(R)\quad$ $\Longleftrightarrow\quad\boldsymbol{D}\otimes_{R}^{\mathbf{L}}\boldsymbol{X}\in\mathcal{B}_{\boldsymbol{D}}(R);$
$\boldsymbol{Y}\in\mathcal{B}_{\boldsymbol{D}}(R)\quad$ $\Longleftrightarrow\quad\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{Y})\in\mathcal{A}_{\boldsymbol{D}}(R);$
$\boldsymbol{X}\in\mathcal{F}(R)\quad$ $\Longleftrightarrow\quad\boldsymbol{D}\otimes_{R}^{\rm\mathbf{L}}\boldsymbol{X}\in\mathcal{I}(R);$
$\boldsymbol{Y}\in\mathcal{I}(R)\quad$ $\Longleftrightarrow\quad\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{Y})\in\mathcal{F}(R).$
In complementation to Theorem 2.2(4), one has the following.
Theorem 2.6 ([7, Proposition 5.9] & [10, Theorem 19.2.34]).
Let $R$ be a local ring with a dualizing complex and let $\boldsymbol{X}\in\mathcal{C}_{(\square)}^{(\rm f)}$. Then $\operatorname{Gid}_{R}(\boldsymbol{X})=\operatorname{Gid}_{\widehat{R}}(\boldsymbol{X}\otimes_{R}\widehat{R})$.
2.12. Stability results
We close this section with the following results, which will be used in the next section.
Theorem 2.7.
Assume $R$ is a local ring, and let $\boldsymbol{X}$ and $\boldsymbol{P}$ two complexes in $\mathcal{C}_{(\square)}^{(\rm f)}(R)$. Suppose that $\boldsymbol{P}$ is non-acylic and that $\operatorname{pd}_{R}(\boldsymbol{P})<\infty$. Then:
  1. (1)
    ([23, Theorems 5.6 & 5.8]) $\operatorname{H-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{P},\boldsymbol{X}))<\infty\Longleftrightarrow\operatorname{H-dim}_{R}(\boldsymbol{X})<\infty$.
  2. (2)
    ([23, Theorems 5.1 & 5.7]) $\operatorname{H-dim}_{R}(\boldsymbol{P}\otimes_{R}^{\mathbf{L}}\boldsymbol{X})<\infty$Longleftrightarrow $\operatorname{H-dim}_{R}(\boldsymbol{X})<\infty$.
Theorem 2.8.
([9, Corollary 6.4]) Assume $R$ is a local ring with a dualizing complex $\boldsymbol{D}$. Then for a complex $\boldsymbol{X}\in\mathcal{C}_{(\square)}^{(\rm f)}(R)$ one has:
$\operatorname{Gid}_{R}(\boldsymbol{X})<\infty\Longleftrightarrow\operatorname{G-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{X},\boldsymbol{D}))<\infty.$
Theorem 2.9.
Assume $R$ is a local ring, and let $\boldsymbol{Y}$ and $\boldsymbol{P}$ be complexes in $\mathcal{C}_{(\square)}^{(\rm f)}(R)$. Suppose that $\boldsymbol{P}$ is non-acyclic and that $\operatorname{pd}_{R}(\boldsymbol{P})<\infty$. In the case $\operatorname{H-dim}_{R}=\operatorname{G-dim}_{R}$, assume further that $R$ has a dualizing complex.
  1. (1)
    If $\operatorname{Hid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{P},\boldsymbol{Y}))<\infty$, then $\operatorname{Hid}_{R}(\boldsymbol{Y})<\infty$.
  2. (2)
    If $\operatorname{Hid}_{R}(\boldsymbol{Y}\otimes_{R}^{\mathbf{L}}\boldsymbol{P})<\infty$, then $\operatorname{Hid}_{R}(\boldsymbol{Y})<\infty$.
Proof.
If $\operatorname{H-dim}=\operatorname{pd}$, it follows from [8, Lemmas 6.2.10 & 6.2.15]. On the other hand, if $\operatorname{H-dim}=\operatorname{G-dim}$, then the theorem follows from [32, Proposition 4.1(b)] and Proposition 2.4(2). ∎
3. Finite homological dimensions of Ext of complexes
In this section, we derive results regarding the finiteness of homological dimension of Ext of complexes, and conclude the corresponding results for Ext of modules. The motivation arises from the [37, Section 6], where the respective authors obtained a result concerning finiteness of the injective dimension of Ext of complexes ([37, Proposition 4.2]) using [37, Lemma 4.1] and Bass Series, and subsequently derived the corresponding result for Ext of modules. Inspired by their approach, we first derive counteparts of the lemma mentioned considering other homological dimensions and then use Poincaré Series or stability results for concluding our results.
Lemma 3.1.
Let
$\boldsymbol{X}:\cdots\longrightarrow X_{l+1}\xrightarrow{\partial_{l+1}}X_{l}\xrightarrow{\partial_{l}}X_{l-1}\rightarrow\cdots$
be a homologically bounded complex. If $\operatorname{Gid}_{R}(\operatorname{H}_{i}(\boldsymbol{X}))<\infty$ for all $i,$ then $\operatorname{Gid}_{R}(\boldsymbol{X})<\infty$.
Proof.
The proof is the same as the proof of [37, Lemma 4.1] but now replacing $\operatorname{Gid}$ by $\operatorname{id}$, considering the complex $\boldsymbol{I}$ of the proof as a resolution of Gorenstein injective modules of $\operatorname{H}_{0}(\boldsymbol{X})$, and using Theorem 2.1(4). ∎
Lemma 3.2.
Let
$\boldsymbol{X}:\cdots\longrightarrow X_{l+1}\xrightarrow{\partial_{l+1}}X_{l}\xrightarrow{\partial_{l}}X_{l-1}\rightarrow\cdots$
be a complex in $\mathcal{C}_{(\square)}(R)$. Suppose further that $\boldsymbol{X}\in\mathcal{C}^{(\rm{f})}(R)$ in the case $\operatorname{H-dim}_{R}=\operatorname{G-dim}$. If $\operatorname{H-dim}_{R}(\operatorname{H}_{i}(\boldsymbol{X}))<\infty$ for all $i,$ then $\operatorname{H-dim}_{R}(\boldsymbol{X})<\infty$.
Proof.
Let $n$ be the number of non-zero integers $i$ such that $\operatorname{H}_{i}(\boldsymbol{X})\not=0$. If $n=0,$ then $\boldsymbol{X}$ is acyclic and hence $\operatorname{H-dim}_{R}(\boldsymbol{X})=-\infty<\infty$. Now, we show by induction that the lemma holds when $n\geq 1$.
If $n=1,$ then there exists an unique integer $i$ such that $\operatorname{H}_{i}(\boldsymbol{X})\not=0$. Without loss of generality, we may assume $i=0$. Set $Z=\operatorname{Coker}(\partial_{1})$. Let
$\boldsymbol{Y}:\cdots\rightarrow 0\rightarrow Z\xrightarrow{\tilde{\partial}_{0}}X_{-1}\xrightarrow{\partial_{-1}}X_{-2}\rightarrow\cdots,$
where ~$\tilde{\partial}_{0}$ is the map induced of $\partial_{0}$. The natural homomorphism $\boldsymbol{X}\to\boldsymbol{Y}$ is a quasi-isomorphism, so $\boldsymbol{X}\simeq\boldsymbol{Y}$ in $\mathcal{D}(R)$. Also, we may note that the natural homomorphism $\operatorname{H}_{0}(\boldsymbol{X})\to\boldsymbol{Y}$, is a quasi-isomorphism, so $\operatorname{H}_{0}(\boldsymbol{X})\simeq\boldsymbol{Y}$ in $\mathcal{D}(R)$. Thus, $\operatorname{H}_{0}(\boldsymbol{X})\simeq\boldsymbol{X}$ in $\mathcal{D}(R)$. As $\operatorname{H-dim}_{R}(\operatorname{H}_{0}(\boldsymbol{X}))<\infty,$ it follows that $\operatorname{H-dim}_{R}(\boldsymbol{X})<\infty$.
Suppose $n>1$. We may suppose that $\sup\{i\in\mathbb{Z}:\operatorname{H}_{i}(\boldsymbol{X})\neq 0\}=0$. As above, if $Z=\operatorname{Coker}(\partial_{1})$ and $\boldsymbol{Y}$ is the complex
$\boldsymbol{Y}:\cdots\rightarrow 0\rightarrow Z\xrightarrow{\tilde{\partial}_{0}}X_{-1}\xrightarrow{\partial_{-1}}X_{-2}\rightarrow\cdots,$
then $\boldsymbol{X}\simeq\boldsymbol{Y}$ in $\mathcal{D}(R)$. Since $\operatorname{H-dim}_{R}(\operatorname{H}_{0}(\boldsymbol{X}))<\infty$, then there exists a bounded complex
$\boldsymbol{T}:\cdots\rightarrow T_{i+1}\rightarrow T_{i}\rightarrow\cdots\rightarrow T_{1}\rightarrow T_{0}\rightarrow 0$
such that $\operatorname{H-dim}_{R}(T_{i})=0$ for all $i\geq 0$, and
$\operatorname{H}_{i}(\boldsymbol{T})=\begin{cases}\operatorname{H}_{0}(\boldsymbol{X}),&\text{if }i=0\\ 0,&\text{if }i>0.\end{cases}$
Let $\boldsymbol{C}$ be the complex
$\boldsymbol{C}:\cdots\rightarrow T_{1}\rightarrow T_{0}\rightarrow Z\xrightarrow{\tilde{\partial}_{0}}X_{-1}\xrightarrow{\partial_{-1}}X_{-2}\rightarrow\cdots,$
where $T_{0}\to Z$ is the composition of the maps $T_{0}\to\operatorname{H}_{0}(X)$ and $\operatorname{H}_{0}(X)\to Z$. Let us consider that the complexes $\boldsymbol{Y}$ and $\boldsymbol{C}$ are indexing such that $Y_{-1}=Z$ and $C_{0}=T_{0}$. Then we have naturally an exact sequence
$0\to\boldsymbol{Y}\to\boldsymbol{C}\to\boldsymbol{T}\to 0$
Note that $\operatorname{H}_{i}(\boldsymbol{C})=0$ for all $i\geq-1$, and $\operatorname{H}_{i}(\boldsymbol{C})=\operatorname{H}_{i+1}(\boldsymbol{X})$ for $i\leq-2$. Hence, the number of non-zero homologies of $\boldsymbol{C}$ is $n-1$. Thus, by induction, $\operatorname{H-dim}_{R}(\boldsymbol{C})<\infty$. Then from Theorem 2.1(1) or (3), we see that $\operatorname{H-dim}_{R}(\boldsymbol{Y})<\infty$. Thus, since $\boldsymbol{X}\simeq\boldsymbol{Y}$, we see that $\operatorname{H-dim}_{R}(\boldsymbol{X})<\infty$. ∎
Theorem 3.3.
Assume $R$ is local, and let $\boldsymbol{M}$ and $\boldsymbol{N}$ be non-acyclic complexes in $\mathcal{C}_{(\square)}^{\rm(f)}(R)$ such that $\operatorname{pd}_{R}(\boldsymbol{M})<\infty$.
  1. (1)
    If $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(\boldsymbol{M},\boldsymbol{N}))<\infty$ for all $i$, then $\operatorname{H-dim}_{R}(\boldsymbol{N})<\infty$.
  2. (2)
    Suppose that $R$ has a dualizing complex. If $\operatorname{Gid}_{R}(\operatorname{Ext}_{R}^{i}(\boldsymbol{M},\boldsymbol{N}))<\infty$ for all $i$, then $\operatorname{Gid}_{R}(\boldsymbol{N})<\infty$.
Proof.
We know that $\operatorname{Ext}_{R}^{i}(\boldsymbol{M},\boldsymbol{N})=\operatorname{H}_{-i}(\mathbf{R}\mathrm{Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))$ for all $i$. From [6, A.4.4 and A.5.4], ones sees that $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N})\mathcal{C}_{(\square)}^{\rm(f)}(R)$. By Lemma 3.2 (resp. Lemma 3.1), we conclude $\operatorname{H-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$ (resp. $\operatorname{Gid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$). It follows from Theorem 2.7(1) (resp. Theorem 2.9(1)) that $\operatorname{H-dim}_{R}(\boldsymbol{N})<\infty$. (resp. $\operatorname{Gid}_{R}(\boldsymbol{N})<\infty$), concluding (1) (resp. (2)). ∎
Corollary 3.4.
Assume $R$ is local, and let $M$ and $N$ be non-zero finitely generated $R$-modules such that $\operatorname{pd}_{R}(M)<\infty$.
  1. (1)
    If $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq\operatorname{pd}_{R}(M)$, then $\operatorname{H-dim}_{R}(N)<\infty$.
  2. (2)
    Suppose that $R$ has a dualizing complex. If $\operatorname{Gid}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq\operatorname{pd}_{R}(M)$, then $\operatorname{Gid}_{R}(N)<\infty$.
Corollary 3.5.
Assume $R$ is local, and let $M$ be a non-zero finitely generated $R$-module such that $\operatorname{pd}_{R}(M)<\infty$ and $\operatorname{Gid}_{R}(\operatorname{Ext}_{R}^{i}(M,M))<\infty$ for all $0\leq i\leq t-\operatorname{depth}(M)$. Then $R$ is Gorenstein.
Proof.
We may assume that $R$ is complete. Then $R$ has a dualizing complex. By Corollary 3.4(2), then $\operatorname{Gid}_{R}(M)<\infty$. Then the result follows from [6, Theorem 6.3.2]. ∎
Item (2) of the next theorem provides a positive answer to the Gorenstein version of [18, Question 4.5].
Theorem 3.6.
Assume $R$ is local, and let $\boldsymbol{M}$ and $\boldsymbol{N}$ be non-acyclic complexes in $\mathcal{C}_{(\square)}^{(\rm f)}(R)$. Suppose that $\operatorname{id}_{R}(\boldsymbol{N})<\infty$. In the case $\operatorname{H-dim}=\operatorname{G-dim}$, assume further that $R$ has a dualizing complex.
  1. (1)
    If $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(\boldsymbol{M},\boldsymbol{N}))<\infty$ for all $i$, then $\operatorname{Hid}_{R}(\boldsymbol{M})<\infty$.
  2. (2)
    If $\operatorname{Gid}_{R}(\operatorname{Ext}_{R}^{i}(\boldsymbol{M},\boldsymbol{N}))<\infty$ for all $i$, then $\operatorname{G-dim}_{R}(\boldsymbol{M})<\infty$.
Proof.
We know that $\operatorname{Ext}_{R}^{i}(\boldsymbol{M},\boldsymbol{N})=\operatorname{H}_{-i}(\mathbf{R}\mathrm{Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))$. Additionally, by [6, A.4.4 and A.5.2], we see that $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N})\in\mathcal{C}_{(\square)}^{(\rm f)}(R)$. In the case (1) (resp. (2)), by assumption, $\operatorname{H-dim}_{R}(\operatorname{H}_{i}(\mathbf{R}\mathrm{Hom}_{R}(\boldsymbol{M},\boldsymbol{N})))<\infty$ (resp. $\operatorname{Gid}_{R}(\operatorname{H}_{i}(\mathbf{R}\mathrm{Hom}_{R}(\boldsymbol{M},\boldsymbol{N})))<\infty$) for all $i$. Thus, by Lemma 3.1 (resp. Lemma 3.2), we have that $\operatorname{H-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$ (resp. $\operatorname{Hid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$) for all $i$.
The conclusion of item (1) for $\operatorname{H-dim}_{R}=\operatorname{pd}_{R}$ it follows from [8, Lemma 6.2.12].
It is remaining to prove (1) for $\operatorname{H-dim}_{R}=\operatorname{G-dim},$ and (2). For proving both cases, either by Theorems 2.2, or by assumption and Theorem 2.6, we can let $\boldsymbol{D}$ be a dualizing complex of $R$. Since $\operatorname{id}_{R}(\boldsymbol{N})<\infty$, then $\boldsymbol{N}\in\mathcal{B}_{\boldsymbol{D}}(R)$ and hence $\operatorname{pd}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{N}))<\infty$ by the Foxby equivalence.
We have
$\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N})$ $\simeq\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{D}\otimes_{R}^{\mathbf{L}}\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{N}))$ because $\text{because }\boldsymbol{N}\in\mathcal{B}_{\boldsymbol{D}}(R)$
[…]
In the case (1) for $\operatorname{H-dim}_{R}=\operatorname{G-dim}_{R}$, we have that $\operatorname{G-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$, concluding from Theorem 2.7(2) that $\operatorname{G-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{D}))<\infty$. Hence $\operatorname{Gid}_{R}(\boldsymbol{M})<\infty$ by Theorem 2.8.
As to (2), we have that $\operatorname{Gid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$. Then from the last isomorphism and Theorem 2.9(2), we conclude that $\operatorname{Gid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{D}))<\infty$. Since $M\simeq\operatorname{\mathbf{R}Hom}_{R}(\operatorname{\mathbf{R}Hom}_{R}(M,\boldsymbol{D}),\boldsymbol{D})$ by [6, A.8.5], then it follows from Theorem 2.8 that $\operatorname{G-dim}_{R}(\boldsymbol{M})<\infty$.
Corollary 3.7.
Assume $R$ is Cohen-Macaulay local, and let $M$ and $N$ be non-zero finitely generated $R$-modules such that $\operatorname{id}_{R}(N)<\infty$. In the case $\operatorname{H-dim}_{R}=\operatorname{G-dim}_{R}$, assume further that $R$ has canonical module. Then
  1. (1)
    If $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t-\operatorname{depth}(M)$, then $\operatorname{Hid}_{R}(M)<\infty$.
  2. (2)
    If $\operatorname{Gid}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t-\operatorname{depth}(M)$, then $\operatorname{G-dim}_{R}(M)<\infty$.
When $\operatorname{H-dim}_{R}=\operatorname{pd}_{R}$, note that Corollary 3.7(1) recovers the celebrated theorem given by Foxby in [15, Corollary 4.4], which states that if there exists a (finitely generated) non-zero $R$-module $M$ with both finite injective dimension and projective dimension, then $R$ is Gorenstein. On the other hand, via Corollary 3.7(2), we obtain the following Gorenstein criterion that generalizes [37, Corollary 4.4] and [18, Corollary 4.7].
Corollary 3.8.
Assume $R$ is Cohen-Macaulay local, and let $M$ be a non-zero finitely generated $R$-module such that $\operatorname{id}_{R}(M)<\infty$ and $\operatorname{Gid}_{R}(\operatorname{Ext}_{R}^{i}(M,M))<\infty$ for all $0\leq i\leq t-\operatorname{depth}(M)$. Then $R$ is Gorenstein.
Proof.
By Corollary 3.7(2), $\operatorname{G-dim}_{R}(M)<\infty$. Then the result follows from [6, Theorem 6.3.2]. ∎
4. Finite Homological dimensions of deficiency modules
Setup.
From now until the end of the text, $R$ is local, and unless otherwise stated, all $R$-modules will be finitely generated.
In this section, we derive some results about the implication of having modules of deficiency with certain finite homological dimensions. For this, we assume here that $R$ has a dualizing complex $\boldsymbol{D}$. Moreover, we suppose that $\boldsymbol{D}$ is normalized in the sense that $\operatorname{Hom}_{R}(k,\boldsymbol{D})\simeq\Sigma^{d}k$, where $d$ is the dimension of $R$.
The existence of $\boldsymbol{D}$ is equivalent to $R$ being a factor of a Gorenstein local ring $S$, i.e., $R$ is an epimorphic image of $S$. In that context, Schenzel [33] generalized the notion of a canonical module of a ring in the following sense: For an $R$-module $M$, set
$K^{i}(M):=\operatorname{Ext}_{S}^{\dim(S)-i}(M,S)$
for all $i\in\mathbb{Z}$. For each $i$, the $R$-module $K^{i}(M)$ is called the $i$-th deficiency module of $M$. The $R$-module $K(M):=K^{\dim_{R}(M)}(M)$ is called the canonical module of $M$. The modules of deficiency satisfy the following isomorphisms:
$H^{i}_{\mathfrak{m}}(M)\cong\text{Hom}_{R}(K^{i}(M),E_{R})\text{ and }K^{i}(M)\cong\operatorname{H}_{i-d}(\operatorname{\mathbf{R}Hom}_{R}(M,\boldsymbol{D})),$
where $E_{R}$ is the hull envelope of the residue field of $R$. Note that the first isomorphism shows that $K^{i}(M)=0$ for $i<0pt(M)$ or $i>\dim(M)$, and $K^{i}(M)\neq 0$ for $i=0pt(M)$ and $i=\dim(M)$. For further details about deficiency modules, we refer the reader to [33, 34].
In [13, 17], the authors investigated the consequences of the finiteness of classical homological dimensions, namely, projective and injective dimensions, of the deficiency modules and the canonical module. Motivated by their work, we apply Theorem 3.6 to extend those results to a broader class of homological dimensions considered in this paper. We also emphasize that our approach differs from theirs in both method and perspective.
Theorem 4.1.
Let $M$ be an $R$-module. Then
  1. (1)
    If $\operatorname{H-dim}_{R}(K^{i}(M))<\infty$ for all $\operatorname{depth}(M)\leq i\leq\dim(M)$, then $\operatorname{Hid}_{R}(M)<\infty$.
  2. (2)
    If $\operatorname{Hid}_{R}(K^{i}(M))<\infty$ for all $\operatorname{depth}(M)\leq i\leq\dim(M)$, then $\operatorname{H-dim}_{R}(M)<\infty$.
Proof.
For each $i$, we note that $K^{i}(M)=\operatorname{H}_{i-d}(\operatorname{\mathbf{R}Hom}_{R}(M,\boldsymbol{D}))=\operatorname{Ext}_{R}^{d-i}(M,\boldsymbol{D})$. Since $K^{i}(M)=0$ for $i<0pt(M)$ or $i>\operatorname{dim}_{R}(M)$, by assumption we see have that all modules $\operatorname{Ext}_{R}^{i}(M,\boldsymbol{D})$ have finite $\operatorname{H-dim}_{R}$ or all modules have finite $\operatorname{Hid}$. Then the result follows from [37, Proposition 4.2] and Theorem 3.6. ∎
The following corollary contains the counterparts of [14, Proposition 3.5(b)] with Gorenstein dimensions.
Corollary 4.2.
Let $M$ be an $R$-module. If $M$ is Cohen-Macaulay, then
  1. (1)
    $\operatorname{H-dim}_{R}(K(M))<\infty$ if and only if $\operatorname{Hid}_{R}(M)<\infty$.
  2. (2)
    $\operatorname{H-dim}_{R}(M)<\infty$ if and only if $\operatorname{Hid}_{R}(K(M))<\infty$.
Proof.
In view of [33, Theorem 1.14], it is enough to prove (1). One implication follows from Theorem 4.1(1). For the other, suppose that $\operatorname{Hid}_{R}(M)<\infty$. Since $M\cong K(K(M))$, then $K(M)$ is Cohen-Macaulay ([33, Theorem 1.14] again) and from Theorem 4.1(2) it follows that $\operatorname{H-dim}_{R}(K(M))<\infty$. ∎
5. On Finite Homological dimension of Ext modules over CM rings
In this section, we establish a duality of certain Ext modules. The following theorem will be often used in this section.
Proposition 5.1 (Proposition 2.4 & [6, Theorem 3.4.11]).
Assume $R$ is Cohen-Macaulay with canonical module $\omega$, and let $M$ be an $R$-module.
  1. (1)
    $\operatorname{G-dim}_{R}(M)<\infty$ if and only if $M\in\mathcal{A}_{\omega}(R)$.
  2. (2)
    $\operatorname{Gid}_{R}(M)<\infty$ if and only if $M\in\mathcal{B}_{\omega}(R)$.
  3. (3)
    $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(\omega,M))<\infty$ if and only if $\operatorname{Hid}_{R}(M)<\infty$.
  4. (4)
    $\operatorname{H-dim}_{R}(M)<\infty$ if and only if $\operatorname{Hid}_{R}(M\otimes_{R}\omega)<\infty$.
In [12, Corollary 42(3)], although the authors provide an independent proof of Proposition 5.1, this proposition assures that their vanishing hypotheses are not necessary.
Corollary 5.2.
Assume $R$ is Cohen-Macaulay with a canonical module $\omega$, and let $N$ be a non-zero $R$-module. Then $R$ is Gorenstein in any of the following cases:
  1. (1)
    $\operatorname{pd}_{R}(N)<\infty$ and $\operatorname{G-dim}_{R}(\operatorname{Hom}_{R}(\omega,N))<\infty$.
  2. (2)
    $\operatorname{G-dim}_{R}(N)<\infty$ and $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(\omega,N))<\infty$.
Proof.
Proposition 5.1 says that in (1), we have $\operatorname{Gid}_{R}(N)<\infty$ while in (2), we have $\operatorname{id}_{R}(N)<\infty$. The result follows from [6, Theorem 6.3.2]. ∎
Proposition 5.3.
Assume $R$ is Cohen-Macaulay with canonical module $\omega$, and let $n$ be a non-negative integer. Let $M$ and $N$ be $R$-modules such that $\operatorname{G-dim}_{R}(N)<\infty$. Then $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(M,N))<\infty$ if and only if $\operatorname{Hid}_{R}(\operatorname{Hom}_{R}(M,N\otimes_{R}\omega))<\infty$. In this case, the natural map
$\Phi_{M,N}:\operatorname{Hom}_{R}(M,N)\otimes_{R}\omega\to\operatorname{Hom}_{R}(M,N\otimes_{R}\omega)$
is an isomorphism.
Proof.
Suppose that $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(M,N))<\infty$. In particular, $\operatorname{G-dim}(\operatorname{Hom}_{R}(M,N))<\infty$. Note that $\operatorname{Hom}_{R}(M,N)$ and $N$ belong to $\mathcal{A}_{\omega}(R)$. Then by [27, Lemma 5.1], the map $\Phi_{M,N}$ is an isomorphism. Hence, it follows from Proposition 5.1(3) that $\operatorname{Hid}_{R}(\operatorname{Hom}_{R}(M,N\otimes_{R}\omega))<\infty$.
Conversely, suppose that $\operatorname{Hid}_{R}(\operatorname{Hom}_{R}(M,N\otimes_{R}\omega))<\infty$. Then by Proposition 5.1(3), $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(\omega,\operatorname{Hom}_{R}(M,N\otimes_{R}\omega))<\infty$. Moreover,
$\operatorname{Hom}_{R}(\omega,\operatorname{Hom}_{R}(M,N\otimes_{R}\omega))$ $\cong\operatorname{Hom}_{R}(\omega\otimes_{R}M,N\otimes_{R}\omega)$
$\cong\operatorname{Hom}_{R}(M,\operatorname{Hom}_{R}(\omega,N\otimes_{R}\omega))$
$\cong\operatorname{Hom}_{R}(M,N),$
where the last isomorphism is because $N\in\mathcal{A}_{\omega}(R).$ Thus, $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(M,N))<\infty$. ∎
Let $\mathcal{A}$ be an abelian category and let $\mathcal{X}$ a class of objects in $\mathcal{A}$. We say that $\mathcal{X}$ is closed under short exact sequences if, in each short exact sequence, the fact that two objects are in $\mathcal{X}$ implies that the third so is. We say that $\mathcal{X}$ is closed under isomorphisms if every object isomorphic to $\mathcal{X}$, is in $\mathcal{X}$.
Lemma 5.4.
Let $\mathcal{X}$ be a class of the category of $R$-modules closed under short exact sequences and isomorphisms. Let $M$ and $N$ be $R$-modules, and $n$ be a non-negative integer. Suppose that $\operatorname{Ext}_{R}^{i}(M,N)$ and $N$ belong to $\mathcal{X}$ for all $0\leq i\leq n$. Then $\operatorname{Hom}_{R}(\Omega^{i}(M),N)\in\mathcal{X}$ for all $0\leq i\leq n$.
Proof.
We must show that for all $0\leq j\leq n$, we have $\operatorname{Hom}_{R}(\Omega^{j}(M),N))\in\mathcal{X}$. We prove this by induction on $j$. For $j=0$, this holds by hypothesis. Now suppose $0<j\leq n$ and that $\operatorname{Hom}_{R}(\Omega^{j-1}(M),N))<\infty$. We have an exact sequence
[…]
which yields
[…]
since $\operatorname{Ext}^{1}_{R}(\Omega^{j-1}(M),N)\simeq\operatorname{Ext}^{j}_{R}(M,N)$. Such an exact sequence splits into
[…]
and
[…]
Since $\mathcal{X}$ is closed under short exact sequences and isomorphisms, it follows that $\operatorname{Hom}_{R}(\Omega^{j}(M),N))\in\mathcal{X}$. ∎
The following theorem improves [27, Lemma 5.1].
Theorem 5.5.
Let $C$ be a semidualizing $R$-module, and $n$ be a non-negative integer. Let $M$ and $N$ be $R$-modules such that $N\in\mathcal{A}_{C}(R)$ and $\operatorname{Ext}_{R}^{i}(M,N)\in\mathcal{A}_{C}(R)$ for all $0\leq i\leq n$. Then, for all $0\leq i\leq n$, there is a natural isomorphism
$\operatorname{Ext}_{R}^{i}(M,N)\otimes_{R}C\cong\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}C).$
Proof.
For $i=0$, the isomorphism holds from [27, Lemma 5.1]. For $i>0$, consider the short exact sequence
[…]
Such an exact sequence yields the following commutative diagram
[…]
where the second row is exact. As to the first one, since $\operatorname{Ext}^{1}_{R}(\Omega^{i-1}(M),N)\cong\operatorname{Ext}^{i}_{R}(M,N)$, by assumption, [35, 1.9(a)] and Lemma 5.4, all $R$-modules of this exact sequence belong to $\mathcal{A}_{C}(R)$. Therefore, the first row turns out to be exact as well by [31, Lemma 3.1.12].
Additionally, by [27, Lemma 5.1], the vertical maps are isomorphisms. Then, from the five lemma, we assure the existence of a natural isomorphism $\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N)\otimes_{R}C\cong\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N\otimes_{R}C)$, from which we conclude the result. ∎
Theorem 5.6.
Assume $R$ Cohen-Macaulay with canonical module $\omega$, and let $n$ be a non-negative integer. Let $M$ and $N$ be $R$-modules such that $\operatorname{H-dim}_{R}(N)<\infty$. Then the following conditions are equivalents:
  1. (1)
    $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq n$.
  2. (2)
    $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega))<\infty$ for all $0\leq i\leq n$.
In any case, we have natural isomorphisms
$\operatorname{Ext}_{R}^{i}(M,N)\otimes_{R}\omega\cong\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega)$
for all $0\leq i\leq n$.
Proof.
We first prove the equivalence. Suppose that $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ (resp. $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega))<\infty$) for all $0\leq i\leq n$. Given $0\leq i\leq n$, we must show that $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega))<\infty$ (resp. $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$). The case $i=0$ follows from Proposition 5.3. Now, suppose $i>0$. Consider the exact sequence
$0\stackrel{{\scriptstyle~{}}}{{\longrightarrow}}\Omega^{i}(M)\stackrel{{\scriptstyle~{}}}{{\longrightarrow}}R^{r_{i}}\stackrel{{\scriptstyle~{}}}{{\longrightarrow}}\Omega^{i-1}(M)\longrightarrow 0$
Then it induces an exact sequence
(5.1) $\scriptstyle{0\longrightarrow\operatorname{Hom}_{R}(\Omega^{i-1}(M),N\otimes_{R}\omega)\longrightarrow\operatorname{Hom}_{R}(R^{r_{i}},N\otimes_{R}\omega)\longrightarrow\operatorname{Hom}_{R}(\Omega^{i}(M),N\otimes_{R}\omega)\longrightarrow\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N\otimes_{R}\omega)\longrightarrow 0}.$
(resp.
(5.2) $\scriptstyle{0\longrightarrow\operatorname{Hom}_{R}(\Omega^{i-1}(M),N)\longrightarrow\operatorname{Hom}_{R}(R^{r_{i}},N)\longrightarrow\operatorname{Hom}_{R}(\Omega^{i}(M),N)\longrightarrow\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N)\longrightarrow 0}\mbox{)}.$
By Proposition 5.1(4), $\operatorname{Hid}_{R}(N\otimes_{R}\omega)<\infty$. Additionally, by Lemma 5.4 and Theorem 2.1, $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(\Omega^{j}(M),N))<\infty$ (resp. $\operatorname{Hid}_{R}(\operatorname{Hom}_{R}(\Omega^{j}(M),N\otimes_{R}\omega))<\infty$) for $j=i-1,i$, so by Proposition 5.3, $\operatorname{Hid}_{R}(\operatorname{Hom}_{R}(\Omega^{j}(M),N\otimes_{R}\omega))<\infty$ (resp. $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(\Omega^{j}(M),N))<\infty$) for $j=i-1,i$. Thus, the first three terms of the exact sequence (5.1) (resp. (5.2)) have finite $\operatorname{Hid},$ (resp. $\operatorname{H-dim}_{R}$) and therefore $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N\otimes_{R}\omega))<\infty$ (resp. $\operatorname{H-dim}_{R}\operatorname{Ext}^{1}_{R}(\Omega^{i-1}(M),N)<\infty$). The result follows from the fact that $\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega)\cong\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N\otimes_{R}\omega)$.
Now, we prove the last statement. Suppose that (1) holds. Then the $R$-modules $N$ and $\operatorname{Ext}_{R}^{i}(M,N)$, for all $i=0,\ldots,n$, have finite Gorenstein dimension. Thus, these modules belong to $\mathcal{A}_{\omega}(R)$. Hence, the isomorphism required follows from Theorem 5.5. ∎
Corollary 5.7.
Assume $R$ is Cohen-Macaulay, and let $M$ and $N$ be non-zero $R$-modules such that $\operatorname{pd}_{R}(N)<\infty$. If $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t-\operatorname{depth}(M)$, then $\operatorname{H-dim}_{R}(M)<\infty$.
Proof.
We may assume that $R$ is complete, and hence $R$ has a canonical module $\omega$. Since $\operatorname{pd}_{R}(N)<\infty$, by Proposition 5.1, we have $\operatorname{id}_{R}(N\otimes_{R}\omega)<\infty$ and, by Theorem 5.6, $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega))<\infty$ for all $0\leq i\leq t-0pt(M)$. Then it follows from Corollary 3.7 that $\operatorname{H-dim}_{R}(M)<\infty$. ∎
The following corollary contains a generalization of [12, Corollary 30] and [28, Theorem 4.2$(3)\Rightarrow(2)$].
Corollary 5.8.
Assume $R$ is Cohen-Macaulay, and let $M$ and $N$ be non-zero $R$-modules such that $\operatorname{pd}_{R}(N)<\infty$ and $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(M,N))<\infty$. If $\operatorname{Ext}_{R}^{i}(M,N)=0$ for all $1\leq i\leq t-0pt(M)$, then $\operatorname{H-dim}_{R}(M)=0$.
Proof.
By Corollary 5.7, we have that $\operatorname{H-dim}_{R}(M)<\infty$. As $\operatorname{pd}_{R}(N)<\infty$, we have that $\operatorname{Ext}_{R}^{i}(M,N)=0$ for all $i>t-0ptM$ by [1, Theorem 4.13], and that $\operatorname{H-dim}_{R}(M)=\sup\{i\geq 0:\operatorname{Ext}_{R}^{i}(M,N)\not=0\}$ by [30, Theorem 2.6(ii)]. Therefore, $\operatorname{H-dim}_{R}(M)=0$. ∎
Combining Proposition 3.4 with Corollary 5.7, we have the following byproduct.
Corollary 5.9.
Assume $R$ is Cohen-Macaulay, and let $M$ and $N$ be non-zero $R$-modules. Suppose that $\operatorname{pd}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t-0pt(M)$. Then $\operatorname{pd}_{R}(M)<\infty$ if and only if $\operatorname{pd}_{R}(N)<\infty$.
Next, we explore the consequences of Theorem 5.6 when in such theorem we consider $N=R$. Recall that the type of an $R$-module $M$ is defined to be the dimension of the $k$-vector space $\operatorname{Ext}^{r}_{R}(k,M)$, where $r=0pt(M)$, and denoted by $\operatorname{r}(M)$.
Corollary 5.10.
Assume $R$ is Cohen-Macaulay with a canonical module $\omega$, and let $n$ be a non-negative integer. Let $M$ be an $R$-module such that $\operatorname{G-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $0\leq i\leq n$. Then for each $0\leq i\leq n$, we have that $\operatorname{Ext}_{R}^{i}(M,R)\otimes_{R}\omega\cong\operatorname{Ext}_{R}^{i}(M,\omega)$, $\operatorname{Tor}^{R}_{1}(\operatorname{Tr}_{R}(\Omega^{i}(M)),\omega)=0$ and $\mu(\operatorname{Ext}^{i}_{R}(M,\omega))=\mu(\operatorname{Ext}^{i}_{R}(M,R))\operatorname{r}(R)$.
Proof.
The isomorphisms follow immediately from Theorem 5.6. Since such isomorphisms are induced by the natural map $\Phi_{M,N}$ in Proposition 5.3, it follows from [1, Theorem 2.8(b)] that $\operatorname{Tor}^{R}_{1}(\operatorname{Tr}_{R}(\Omega^{i}(M)),\omega)=0$ for all $0\leq i\leq n$.
Now, by taking the tensor product by the residue field of $R$ and dimensions, the required equality holds from [5, Proposition 3.3.11(c)]. ∎
Corollary 5.11.
Assume $R$ is Cohen-Macaulay, and let $M$ be an $R$-module of depth $r$ and dimension $s$. Suppose that $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $t-s\leq i\leq t-r$ Then:
  1. (1)
    $\operatorname{H-dim}_{R}(M)<\infty$.
  2. (2)
    $\operatorname{dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))\leq t-i$ for all $i\neq t-s$ and $\dim_{R}(\operatorname{Ext}^{t-s}_{R}(M,R))=s$.
  3. (3)
    $\operatorname{r}(M)=\mu(\operatorname{Ext}^{t-r}_{R}(M,R))\operatorname{r}(R).$
  4. (4)
    If $M$ has type one, then $R$ is Gorenstein.
Proof.
We may assume that $R$ is complete and therefore $R$ admits a canonical module $\omega$. By [5, Corollary 2.1.4], $\operatorname{grade}(M)=t-s$, so $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $i<t-s$ and $\operatorname{Ext}_{R}^{t-s}(M,R)\not=0$. Then by Corollary 5.7, $\operatorname{H-dim}_{R}(M)<\infty$ and (1) it is proved.
On the other hand, since $\operatorname{Ext}_{R}^{i}(M,R)=0$ for $i>t-r$, one can see by assumption that $\operatorname{G-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $i\geq 0$. Thus, given $i\geq 0$, by Corollary 5.10,
$\operatorname{Ext}_{R}^{i}(M,R)\otimes_{R}\omega\cong\operatorname{Ext}_{R}^{i}(M,\omega).$
It follows that the modules $\operatorname{Ext}^{i}_{R}(M,R)$ and $\operatorname{Ext}^{i}_{R}(M,\omega)$ have the same support and so $\dim_{R}(\operatorname{Ext}_{R}^{i}(M,R))=\dim_{R}(\operatorname{Ext}_{R}^{i}(M,\omega))$. In this way, item (2) follows from [33, Lemma 1.9 a) and c)].
Item (3) follows from Corollary 5.10 and [13, Theorem 3.1], and item (4) follows from (3) and [5, Theorem 3.2.10(b)]. ∎
Corollary 5.12.
Assume $R$ is Cohen-Macaulay, and let $M$ be an $R$-module of grade $g$. If $M$ is Cohen-Macaulay and $\operatorname{H-dim}(\operatorname{Ext}_{R}^{g}(M,R))<\infty,$ then $M$ is $\operatorname{H}$-perfect.
Proof.
By [5, Corollary 2.1.4], $g=t-s$ with $s=\dim_{R}(M)$, so by assumption $\operatorname{G-dim}_{R}(\operatorname{Ext}_{R}^{t-s}(M,R))<\infty$. Since $M$ is Cohen-Macaulay, it follows from Corollary 5.11(1), that $\operatorname{H-dim}_{R}(M)<\infty$. It follows from the formula Proposition 2.3(2) that $\operatorname{H-dim}_{R}(M)=t-s=g$, i.e., $M$ is $\operatorname{H}$-perfect. ∎
Remark 5.13.
In [16, Theorem 2 and 2’], Foxby showed that if $R$ is Cohen-Macaulay, then an $R$-module $M$ of grade $g$ is $\operatorname{H}$-perfect if and only if $M$ is Cohen-Macaulay and $\operatorname{Ext}_{R}^{g}(M,R)$ is $\operatorname{H}$-perfect. Note that Corollary 5.12 generalizes the converse of that result.
In view of [18, Theorem 3.3], it should be noted that Corollary 5.11(3) generalizes the Ghosh-Puthenpurakal type formula [18, Theorem 3.5] for Cohen-Macaulay rings and answers negatively [18, Question 3.9] when we are over Gorenstein rings.
Corollary 5.14.
Assume $R$ is Gorenstein. For any $R$-module $M$ one has
$\operatorname{r}(M)=\mu(\operatorname{Ext}_{R}^{\operatorname{G-dim}_{R}(M)}(M,R)).$
Although most of the next result is known, our results (including Theorem 5.6) show that such different characterizations of the Gorenstein property of a Cohen-Macaulay ring come from the same place.
Corollary 5.15.
Assume $R$ is Cohen-Macaulay and let $\omega$ be its canonical module. We let $\_^{\dagger}$ stands for $\operatorname{Hom}_{R}(\_,\omega)$. The following are equivalent:
  1. (1)
    $R$ is Gorenstein.
  2. (2)
    $\operatorname{Gid}_{R}(k)<\infty$.
  3. (3)
    $\operatorname{Gid}(R)<\infty$.
  4. (4)
    $\operatorname{G-dim}_{R}(\omega)<\infty$.
  5. (5)
    $\operatorname{G-dim}_{R}(\omega^{*})<\infty$.
    Now consider the following statements provided that $\operatorname{Ext}^{i}_{R}(\omega,R)=0$ for all $1\leq i\leq t$.
  6. (6)
    $\operatorname{Gid}_{R}((\omega^{*})^{\dagger})<\infty$.
  7. (7)
    $\operatorname{Gid}_{R}(\omega^{*})<\infty$.
  8. (8)
    $\operatorname{G-dim}_{R}((\omega^{*})^{\dagger})<\infty$.
Then $\rm(1)-(5)$ are mutually equivalent, and each of them is equivalent to $\rm(6)$. In addition, $(6)\Rightarrow(7)\Leftrightarrow(8)$.
Proof.
The equivalences $(1)\Leftrightarrow(2)$ and $(3)\Leftrightarrow(4)$ follow directly from Corollary 4.2. Also, Theorem 5.6 asserts immediately that $(3)\Leftrightarrow(5)$. The equivalence $(1)\Leftrightarrow(3)$ is proved in [22, Theorem 2.1].
Now, note that from the vanishing hypothesis and [21, Theorem 2.1(i)] we see that $\omega^{*}$ is maximal Cohen-Macaulay. Therefore $(5)\Leftrightarrow(6)$ and $(7)\Leftrightarrow(8)$ follow from Corollary 4.2. Since $(1)\Leftrightarrow(6)$, it is clear that $(6)\Rightarrow(7)$. ∎
6. Some questions
In this section, we discuss some natural questions motivated by the results established earlier. Our primary question stems from the challenge of extending Corollaries 5.7 and 5.8 to the non-Cohen-Macaulay setting, removing the requirement that $R$ be a Cohen-Macaulay ring.
Question 6.1.
Let $M$ and $N$ be non-zero $R$-modules such that $\operatorname{pd}_{R}(N)<\infty$. Do the following facts hold?
  1. (1)
    If $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t$, then $\operatorname{H-dim}_{R}(M)<\infty$.
  2. (2)
    If $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(M,N))<\infty$ and $\operatorname{Ext}_{R}^{i}(M,N)=0$ for all $1\leq i\leq t$, then $\operatorname{H-dim}_{R}(M)=0$.
The following proposition partially answers Question 6.1 under the assumption that $\operatorname{H\text{-}dim}_{R}=\operatorname{pd}_{R}$.
Proposition 6.2.
Let $M$ and $N$ be non-zero $R$-modules. Let $s$ be a non-negative integer such that $\operatorname{pd}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq s$. Then the equivalence
$\operatorname{pd}_{R}(M)<\infty\Longleftrightarrow\operatorname{pd}_{R}(N)<\infty$
holds in each one of the following cases:
  1. (1)
    If $s=t$ and $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $t+1\leq i\leq 2t+1$.
  2. (2)
    $M$ has finite Gorenstein dimension and $s=t-0pt(M)$
Proof.
In both cases, the implication ()($\Rightarrow$)( ⇒ ) follows from Proposition 3.4. Thus, it remains to prove the implication ()($\Leftarrow$)( ⇐ ).
(1) Suppose that $\operatorname{pd}_{R}(N)<\infty$. By Theorem 2.1 and Lemma 5.4, $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(\Omega^{t}(M),N))<\infty$. Moreover, the vanishing condition shows that $\operatorname{Ext}_{R}^{i}(\Omega^{t}(M),R)=0$ for all $1\leq i\leq t$. Therefore, by [12, Proposition 19], $\operatorname{pd}_{R}(\Omega^{t}(M)^{\ast})<\infty$. Then by [12, Proposition 13], $\Omega^{t}(M)$ is free. Thus, $\operatorname{pd}_{R}(M)<\infty$.
(2) Since $\operatorname{pd}_{R}(N)<\infty$, then $\operatorname{Ext}_{R}^{i}(M,N)=0$ for all $i>\operatorname{G-dim}_{R}(M)=0pt(R)-0pt(M)$. Then by assumption, we can see that $\operatorname{pd}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t$. Therefore, the assertion follows from item (1). ∎
Remark 6.3.
It should be noticed that in Proposition 6.2 the implication $\Rightarrow$ only requires that $\operatorname{pd}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for $0\leq i\leq t-\operatorname{depth}(M)$.
For $N=R$, an answer to Question 6.1(2) is given by [12, Proposition 13]. Next, we present an answer to Question 6.1(1) in the following proposition, which follows from and strengthens [12, Proposition 13] for projective dimension and Gorenstein dimension. For that, we say that an $R$-module $M$ satisfies $(\widetilde{S}_{n})$ if $\operatorname{depth}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})\geq\min\{n,0pt_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\}$ for all prime ideals $\mathfrak{p}$ of $R$.
Proposition 6.4.
Let $M$ be an $R$-module satisfying $(\widetilde{S}_{n})$ with $n\leq t$ a non-negative integer. Suppose that $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $0\leq i\leq t-n$. Then $\operatorname{H-dim}_{R}(M)\leq t-n$.
Proof.
Let $s=t-n$. Then by Theorem 2.1 and Lemma 5.4, $\operatorname{H-dim}_{R}(\Omega^{s}(M)^{\ast})<\infty$. Since $\operatorname{H-dim}_{R}(M^{\ast})<\infty$ and $M$ satisfies $(\widetilde{S}_{n})$, an application of the depth lemma shows that $\Omega^{s}(M)$ satisfies $(\widetilde{S}_{t})$. It follows from [12, Proposition 13] that $\operatorname{H-dim}_{R}(\Omega^{s}(M))=0$. Hence, by Proposition 2.3(3), $\operatorname{H-dim}_{R}(M)\leq s=t-n$. ∎
Remark 6.5.
It is important to mention Proposition 6.4 is significant only when $\operatorname{grade}(M)=0$ or $n=0$. The reason of this is that there are no non-zero $R$-modules $N$ with positive grade satisfying $(\widetilde{S}_{1})$. The proof of this fact is as follows. Suppose that $N$ satisfies $(\widetilde{S}_{1})$ and that $N^{\ast}=0$. Then, $\operatorname{pd}_{R}(\operatorname{Tr}_{R}(N))\leq 1$. As $N$ satisfies $(\widetilde{S}_{1})$, by [36, Theorem 5.8], we have that $\operatorname{Ext}_{R}^{1}(\operatorname{Tr}_{R}(N),R)=0$. Thus, $\operatorname{Tr}_{R}(N)$ is free and hence $N$ as well. But since $N^{\ast}=0,$ then $N=0$.
Corollary 6.6.
Let $M$ be an $R$-module. If $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $0\leq i\leq t$, then $\operatorname{H-dim}_{R}(M)<\infty$.
Note that Proposition 6.4 and Corollary 6.6 give some answers to [28, Question 4.11]. As an application of Corollary 6.6, we can improve [19, Proposition 3.3] as follows:
Corollary 6.7.
Let $M$ be a non-zero $R$-module. Assume further that there exists a non-negative integer $h$ such that $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{h}(M,R))<\infty$ and $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $i\in\{0,\ldots,t\}-\{h\}$, then $\operatorname{H-dim}_{R}(M)=h$.
Proof.
From the assumption we see that $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $0\leq i\leq t$. Then by Corollary 6.6, $\operatorname{H-dim}_{R}(M)<\infty$. Since $\operatorname{H-dim}_{R}(M)=\sup\{i\geq 0:\operatorname{Ext}_{R}^{i}(M,R)\not=0\}$ (see, for instance [26, Lemma 2.3(c)] together with [2, Theorem 1.4]) and $\operatorname{H-dim}_{R}(M)\leq t$, we see from the vanishing condition that $\operatorname{H-dim}_{R}(M)=h$. ∎
The reader can note that Corollaries 5.7 and 5.8, which motivated Question 6.1, were obtained using Theorem 5.6, which was derived from Theorem 5.5. Thus, in view of the Foxby equivalence ([6, Theorem 3.3.2]), trying to answer these questions, we pose the following question:
Question 6.8.
It is possible to obtain a variation of Proposition 5.3 or Theorem 5.5 involving a dualizing complex $\boldsymbol{D}$ instead of $\omega$ or $C$?
Now, motivated by Corollary 5.9, we pose the following:
Question 6.9.
Let $M$ and $N$ be non-zero $R$-modules such that $\operatorname{pd}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t$. Then does the equivalence
$\operatorname{id}_{R}(M)<\infty\Longleftrightarrow\operatorname{id}_{R}(N)<\infty$
hold?
One can note that the implication ()($\Leftarrow$)( ⇐ ) holds by Theorem 3.6. On the other hand, the other implication has a positive answer if $M$ is maximal Cohen-Macaulay. Indeed, by the Bass conjecture, $R$ is Cohen-Macaulay, and we may assume (passing to completion if necessary) that $R$ admits a canonical module. Then $M\cong\omega^{n}$ for some $n>0$ and $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(\omega,N))<\infty$, obtaining thus the desired conclusion from Proposition 5.1(3).
Based on the blend between the Gorenstein dimensions, the canonical module, and its duals in Corollary 5.15, we close the paper by asking the following.
Question 6.10.
Assume $R$ is Cohen-Macaulay with canonical module $\omega$, and let $\_^{\dagger}$ stands for $\operatorname{Hom}_{R}(\_,\omega)$. Does any of the conditions
  1. (1)
    $\operatorname{Gid}_{R}(\omega^{*})<\infty$;
  2. (2)
    $\operatorname{G-dim}_{R}((\omega^{*})^{\dagger})<\infty$.
imply that $R$ must be Gorenstein?
Acknowledgment. This work was completed while the third author was visiting the University of Nebraska-Lincoln. He is grateful for the hospitality. He also thanks Roger and Sylvia Wiegand, Ryo Takahashi, Hailong Dao, Paulo Martins, Nawaj KC, Zachary Nason, Anna Brosowsky, Andrew Soto Levins and Kesavan Mohana Sundaram for some discussions about this work. In addition, he also thanks to Lars W. Christensen and Henrik Holm for their response to a question about completion and Gorenstein injective dimension. Additionally, the authors are grateful to Ryo Takahashi, Hailong Dao and Nawaj KC for suggesting some examples, and to Dipankar Ghosh for correcting some typos.

### v2

@@
 In this direction, we establish several finiteness criteria for homological dimensions.
-Our results include the consequences of the finiteness of the Gorenstein (injective) dimension of the deficiency modules as well as a duality for certain Ext modules, all of finite (Gorenstein) projective dimension.
+Under such finiteness conditions, our main result is a new duality for certain Ext modules when tensored by canonical modules.
 Key words and phrases:
@@
 After the astonishing ideas of Auslander, Buchsbaum, and Serre culminating in the characterization of the regular property of a Noetherian local ring by the finiteness of the projective dimension of its residual field, homological algebra has played a fundamental role in commutative ring theory.
-Indeed, this foundation enabled diverse characterizations of ring-theoretic properties via finiteness conditions on homological dimensions (e.g., projective, complete intersection, Gorenstein dimensions, and their injective versions), see for instance some classical references [1, 2, 6, 15, 29].
-Establishing criteria for the finiteness of homological dimensions as well as their prescription – especially via Ext vanishing – is a very active topic in homological commutative algebra, containing a vast literature [1, 4, 6, 9, 12, 13, 18, 19, 21, 23, 27, 28, 26, 37, 36].
-Recently, an approach to study the implications of vanishing of Ext is under the condition that certain Hom module has some finite homological dimension, as evidenced in [11, 12, 20, 28].
+Indeed, this foundation enabled diverse characterizations of ring-theoretic properties via finiteness conditions on homological dimensions (e.g., projective, complete intersection, Gorenstein dimensions, and their injective versions), see, for instance, some classical references [1, 2, 6, 14, 25].
+Establishing criteria for the finiteness of homological dimensions as well as their prescription – especially via vanishing of Ext modules – is a very active topic in homological commutative algebra, containing a vast literature including but not limited to [1, 4, 6, 9, 11, 12, 15, 16, 18, 19, 23, 24, 22, 32, 31].
+Recently, an approach to study the implications of vanishing of Ext is under the condition that certain Hom module has some finite homological dimension, as evidenced in [11, 17, 24].
 In addition, Ghosh-Puthenpurakal [15] and Zargar-Gheibi [32] studied the consequences of having Ext modules with finite injective dimension, deriving criteria for finiteness for the projective dimension of modules.
@@
 Thus, the standard assumption in the present text does not revolve around the usual Ext vanishing hypothesis but rather the finiteness of its homological dimensions.
-Zargar and Gheibi proved in [37] that for non-zero finitely generated modules $M$ and $N$ over a local ring $R$ such that $\operatorname{id}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $i$, the finiteness of the projective dimension of $M$ is equivalent to that of injective dimension of $N$.
-Then, what can one tell if one consider the Ext modules $\operatorname{Ext}_{R}^{i\geq 0}(M,N)$ with another finite homological dimension under the additional condition that $\operatorname{pd}_{R}(M)<\infty$ or $\operatorname{id}_{R}(N)<\infty$?
-Based on their approach employed in [37, Section 4], we obtain results concerning the finiteness of the homological dimension of Ext of complexes (see Theorem 3.3 and Theorem 3.6), obtaining consequently the following:
+Zargar and Gheibi proved in [32] that for non-zero finitely generated modules $M$ and $N$ over a Noetherian local ring $R$ such that $\operatorname{id}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $i$, the finiteness of the projective dimension of $M$ is equivalent to that of injective dimension of $N$.
+Then, what can one tell if one considers the Ext modules $\operatorname{Ext}_{R}^{i\geq 0}(M,N)$ with another finite homological dimension under the additional condition that $\operatorname{pd}_{R}(M)<\infty$ or $\operatorname{id}_{R}(N)<\infty$?
+We address this question in the following theorem.
 Theorem 1.1 (Corollary 2.3 & Corollary 2.6).
-Let $R$ be a local ring of dimension $d$, and let $M$ and $N$ be non-zero finitely generated $R$-modules.
+Let $R$ be a Noetherian local ring of dimension $d$, and let $M$ and $N$ be non-zero finitely generated $R$-modules.
 Let $\mathcal{H}$ be any of the homological dimensions $\operatorname{pd},\operatorname{G-dim},\operatorname{Gid}$, and suppose that $\mathcal{H}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq d-\operatorname{depth}(M)$.
@@
 Theorem 1.1 not only gives criteria for detecting the finiteness of homological dimension of modules but also allows one to obtain Gorenstein criteria as verified in Corollary 2.4 and Corollary 2.7.
-In addition, it is worth noting that Theorem 1.1 says that the counterparts with Gorenstein dimensions of the results of Foxby in [14, Proposition 3.5(b)] are true.
-Furthermore, it is worth to note that item (2.a) allows to recover the celebrated result of Foxby given in [15, Corollary 4.4], and that item (2.c) gives a positive answer to the Gorenstein version of the question in [18, Question 4.5].
-In our study, we also consider the deficiency modules, which were introduced by Schenzel [33].
-They are an example of Ext modules, so what can one tell if all of the deficiency modules of a module have some finite homological dimension?
-Fiel and the first author [13] proved that the finiteness of the projective (resp.
-injective) dimension of all the deficiency modules of a finitely generated module $M$ over a local ring $R$ implies that $M$ has finite injective (resp.
+In addition, it is worth noting that Theorem 1.1 says that the counterparts with Gorenstein dimensions of the results of Foxby in [13, Proposition 3.5(b)] are true.
+Furthermore, item (2.a) allows us to recover a celebrated result of Foxby given in [14, Corollary 4.4], and item (2.c) gives a positive answer to the Gorenstein version of the question in [15, Question 4.5].
+In our study, we also consider deficiency modules, which were introduced by Schenzel [29].
+Fiel and Holanda [12] proved that the finiteness of the projective (resp.
+injective) dimension of all deficiency modules of a finitely generated module $M$ over a local ring $R$ implies that $M$ has finite injective (resp.
 projective) dimension.
-We not only prove this result with a different approach, but also that its Gorenstein version is true.
+We not only prove this result with a different approach, but also, we establish its Gorenstein version.
+The converse is assured for Cohen-Macaulay modules, extending Foxby [13].
 Theorem 1.2 (Theorem 3.1).
-Let $R$ be a local ring with a dualizing complex and let $M$ be a finitely generated $R$-module.
+Let $R$ be a Noetherian local ring with a dualizing complex and let $M$ be a finitely generated $R$-module.
   1. (1)
@@
     If $\operatorname{Gid}_{R}(K^{i}(M))<\infty$ for all $i\geq 0$, then $\operatorname{G-dim}_{R}(M)<\infty$.
-Over Cohen-Macaulay local rings admitting canonical module, we furnish a duality between certain Ext modules under the finiteness of some homological dimensions.
+Our main result is a new duality concerning the canonical module when tensored by certain Ext modules.
 Such a duality is also established for semidualizing modules (Theorem 4.5) and is applied in many directions throughout the paper.
 Theorem 1.3 (Theorem 4.6).
-Assume $R$ Cohen-Macaulay local with canonical module $\omega$, and let $n$ be a non-negative integer.
-Let $\operatorname{H-dim}_{R}$ denote either the projective dimension or the Gorenstein dimension, and $\operatorname{Hid}_{R}$ the injective counterpart.
+Assume $R$ is a Cohen-Macaulay local with canonical module $\omega$, and let $n$ be a non-negative integer.
+Let $\operatorname{H-dim}_{R}$ denote either $\operatorname{pd}$ or $\operatorname{G-dim}$, and $\operatorname{Hid}_{R}$ the injective counterpart.
 Let $M$ and $N$ be finitely generated $R$-modules such that $\operatorname{H-dim}_{R}(N)<\infty$.
@@
 for all $0\leq i\leq n$.
-This result leads to new finiteness criteria of homological dimension, for instance, we prove that if $R$ is a Cohen-Macaulay local ring of depth $t$ and, $M$ and $N$ are non-zero finitely generated $R$-modules such that $\operatorname{pd}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t-\operatorname{depth}(M)$, then finiteness of the projective dimension $M$ is equivalent to the finiteness to that of $N$, see Corollary 5.9.
-Moreover, our theorem allows us to show that over Cohen-Macaulay local rings, the Ghosh-Puthenpurakal type formula [18, Theorem 3.5] is true in other classes of modules than G-perfect modules, and, consequently, furnishes a Gorensteinness criterion for such a ring, see Corollary 5.11.
+Using the behavior of the canonical module described in Theorem 1.3, we derive the following:
+Corollary 1.4 (Corollary 4.8).
+Let $R$ be a Cohen-Macaulay local ring of depth $t$ and, $M$ and $N$ be non-zero finitely generated $R$-modules such that $\operatorname{pd}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t-\operatorname{depth}(M)$, then $\operatorname{pd}_{R}(M)<\infty$ if and only if $\operatorname{pd}_{R}(N)<\infty$.
 The organization of this paper is as follows.
-In Section 2, since our study considers tools from the derived category, we give preliminaries about it.
-In Section 3, we establish results concerning Ext of complexes of finite homological dimension, prove Theorem 1.1, and derive Gorenstein criteria.
+In Section 2, we establish results concerning Ext modules of complexes of finite homological dimension, and prove Theorem 1.1, and derive Gorenstein criteria.
 In Section 3, we study the implications of having deficiency modules with some finite homological dimension.
-In Section 5, we focus our study on modules over Cohen-Macaulay local rings, prove Theorem 5.6, and derive several consequences.
+In Section 4, we prove our main result, namely Theorem 1.3, and derive several consequences.
 In Section 5, we discuss some questions motivated by results in Section 4.
-Next, we establish our setup and provide simple examples of $\operatorname{Ext}$ modules with finite homological dimension to motivate our study.
+Finally, in Section 6, we present some examples of Ext modules with finite homological dimension.
+Now, we establish our setup.
 Setup.
 Throughout this paper, $R$ stands for a Noetherian (commutative unitary) ring.
-Whenever we assume that $R$ is local, its maximal ideal will be denoted by $\mathfrak{m}$ and its residue field by $k$; also, its depth is denoted by $t$.
+Whenever we assume that $R$ is local, its maximal ideal will be denoted by $\mathfrak{m}$ and its residue field by $k$.
+The depth of $R$ is denoted by $t$.
 In that case, for a finitely generated $R$-module $M$, its $i$-th syzygy module and its transpose of Auslander are considered defined by a minimal free resolution of $M$, and denoted by $\Omega^{i}(M)$ and $\operatorname{Tr}_{R}(M)$, respectively.
-Also, for an $R$-module $M$, we let $\operatorname{H-dim}_{R}(M)$ stand for $\operatorname{pd}_{R}(M)$ or $\operatorname{G-dim}_{R}(M)$, and $\operatorname{Hid}_{R}(M)$ for the corresponding injective version.
-For any unexplained terminology, we refer to [5] and [10].
-Example 1.4.
-Assume $R$ is a local ring and let $\underline{x}=x_{1},\ldots,x_{n}$ be a regular sequence.
-Let $N$ be a finitely generated $R$-module such that $\underline{x}N=0$.
-Set $M=R/(\underline{x})$.
-Since the minimal free resolution of $M$ is the Koszul complex of the sequence $\underline{x}$, we have
-$\operatorname{Ext}^{i}_{R}(M,N)\cong N^{\binom{n}{i}},\quad\forall\ i=0,\ldots,n.$
-Thus, $\operatorname{H-dim}_{R}(\operatorname{Ext}^{i}_{R}(M,N))<\infty$ for all $0\leq i\leq n$ if and only if $\operatorname{H-dim}_{R}(N)<\infty.$ Similarly, $\operatorname{Hid}_{R}(\operatorname{Ext}^{i}_{R}(M,N))<\infty$ for all $0\leq i\leq n$ if and only if $\operatorname{Hid}_{R}(N)<\infty.$
-Now, set $L=\oplus_{i=1}^{n}R/(x_{1},\ldots,x_{i})$ and let $X$ be a finitely generated $R$-module such that $\underline{x}$ is $X$-regular. For all $i=1,\ldots,n$, by the Koszul duality we must have
-$\operatorname{Ext}^{i}_{R}(L,X)\cong\operatorname{Ext}^{i}_{R}(R/(x_{1},\ldots,x_{i}),X)\cong X/(x_{1},\ldots,x_{i})X$
-Thus $\operatorname{H-dim}_{R}(X)<\infty$ if and only if $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(L,X))<\infty$ for all $0\leq i\leq n$, and in that case, $\operatorname{H-dim}_{R}(\operatorname{Ext}^{i}_{R}(L,R))=i$, by the Auslander-Buchsbaum and Auslander-Bridge formulas.
-Let $M$ be a finitely generated $R$-module.
-The grade of a finitely generated $R$-module $M$ is defined as $\operatorname{grade}_{R}(M)=\inf\{i\geq 0:\operatorname{Ext}_{R}^{i}(M,R)\not=0\}$.
-The module $M$ is said to be $\operatorname{H-}$perfect if $\operatorname{H-dim}_{R}(M)<\infty$ and $\operatorname{grade}(M)=\operatorname{H-dim}_{R}(M)$.
-Example 1.5.
-Let $M_{1},\ldots,M_{n}$ be H-perfect $R$-modules of grade $g_{1},\ldots,g_{n}$, respectively, and set $M=\bigoplus_{i=1}^{n}M_{i}$. Then
-$\operatorname{Ext}^{g}_{R}(M,R)\cong\begin{cases}\operatorname{Ext}^{g_{i}}_{R}(M_{i},R),&\quad g=g_{i}\\ 0,&\quad\mbox{else}\end{cases}$
-and $\operatorname{H-dim}_{R}(\operatorname{Ext}^{g_{i}}_{R}(M,R))=g_{i}$ by [3, Theorem 6.3].
-An $R$-module $M$ is $p$-spherical if $\operatorname{pd}_{R}(M)\leq p$ and $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $11\leq i\leq p-$ 1.
-Example 1.6.
-Based on the proof of [25, Proposition 6] we can construct spherical non-perfect modules as follows.
-We repeat the proof for convenience.
-Let $M$ be a finitely generated $R$-module with positive grade $n$, and let $F_{\bullet}$ be a free resolution of $M$.
-Consider the truncation
-[…]
-Set $N:=\operatorname{coker}(\phi^{*})$. Since $\operatorname{grade}(M)=n>0$, the complex
-[…]
-is a free resolution of $N$ and
-[…]
-is exact.
-This means that $N$ is $n$-spherical as well as a $(n+1)$th syzygy of $M$ (in the sense of [26]), and also $\operatorname{Ext}^{n}_{R}(N,R)\cong M$.
-It follows that whenever $M$ is such that $\operatorname{pd}_{R}(M)>\operatorname{grade}(M)=n>0$, there exists a $n$-spherical non-perfect $R$-module $N$ such that $\operatorname{Ext}^{n}_{R}(N,R)\cong M$.
-In particular, if $\operatorname{H-dim}_{R}(M)<\infty$ (resp.
-$\operatorname{Hid}_{R}(M)<\infty)$ then $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{n}(N,R))<\infty$ (resp.
-$\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{n}(N,R)<\infty$).
-It is noteworthy that the previous procedure works likewise if one considers resolutions by totally reflexive modules instead.
-In this case, $N$ would still be a non-G-perfect module which is a $(n+1)$th syzygy of $M$ (again in the sense of [26]) such that $\operatorname{G-dim}_{R}(N)\leq n$ and $\operatorname{Ext}^{i}_{R}(N,R)=0$ for all $i\neq 0,n$.
-Example 1.7.
-Let $k$ be a field and consider $R=k[[x,y]]/(xy)$. Then $M=R/(x)$ has infinite free resolution
-[…]
-so that $\operatorname{Ext}^{2i+1}_{R}(M,M)\cong M$ for all $i$, which means that $\operatorname{CI-dim}_{R}(\operatorname{Ext}^{2i+1}_{R}(M,M))=0$ and $\operatorname{pd}_{R}(\operatorname{Ext}^{2i+1}_{R}(M,M))=\infty$.
-2. Preliminaries on Derived Categories
-In this section, we give some preliminaries concerning the derived category.
-For unexplained terminology, we refer to [6, Appendix], and for background on derived category theory, we additionally refer to [10].
-2.1. Complexes and morphisms
-All complexes considered in this section are descending. Let
-$\boldsymbol{X}:\cdots\longrightarrow X_{i+1}\stackrel{{\scriptstyle\partial_{i+1}^{\boldsymbol{X}}}}{{\longrightarrow}}X_{i}\stackrel{{\scriptstyle\partial_{i}^{\boldsymbol{X}}}}{{\longrightarrow}}X_{i-1}\longrightarrow\cdots$
-be a complex of $R$-modules.
-If $X_{i}=0$ for all $i\gg 0$ (resp.
-for all $i\ll 0$), then $\boldsymbol{X}$ is said to be bounded to the left (resp.
-bounded to the right).
-If $\boldsymbol{X}$ is a bounded to the right and to the left, it is said to be bounded.
-We say that $\boldsymbol{X}$ is homologically bounded to the left (resp.
-homologically bounded to the right) if $\operatorname{H}_{i}(\boldsymbol{X})=0$ for all $i\gg 0$.
-If $\boldsymbol{X}$ is homologically bounded to the left and to the right, then $\boldsymbol{X}$ is said to be homologically bounded.
-We recall that an morphism $f:\boldsymbol{X}\rightarrow\boldsymbol{Y}$ of complexes is a family $\{f_{i}:X_{i}\to Y_{i}\}_{i\in\mathbb{Z}}$ of homomorphism of $R$-modules such that $\partial_{i}^{\boldsymbol{Y}}\circ f_{i}=f_{i-1}\circ\partial_{i}^{\boldsymbol{X}}$.
-It is natural to define the composition of morphisms of complexes.
-The category of complexes over $R$ and homomorphism of $R$-modules is denoted by $\mathcal{C}(R)$.
-Note that a morphism $f:\boldsymbol{X}\rightarrow\boldsymbol{Y}$ of complexes induces naturally homomorphisms $\operatorname{H}_{i}(f):\operatorname{H}_{i}(\boldsymbol{X})\rightarrow\operatorname{H}_{i}(\boldsymbol{Y})$.
-We say that $f$ is said to be a quasi-isomorphism if all the induced maps $H_{i}(f)$ are isomorphisms.
-[…]
-An $R$-module $M$ is identified naturally as a complex $\boldsymbol{X}$ such that $X_{0}=M$ and $X_{i}=0$ for all $i\neq 0$.
-2.2. Notation of some full subcategories of $\mathcal{C}(R)$
-We are considering the following full subcategories of $\mathcal{C}(R)$:
+We also consider the derived category of $R$.
+Unless, otherwise specified, we adopt the terminology used in [10].
+In particular, we denote the projective and Gorenstein dimensions of an $R$-module $M$ respectively by $\operatorname{pd}_{R}(M)$ and $\operatorname{G-dim}_{R}(M)$, while the injective dimension and Gorenstein injective dimension of $M$ are respectively denoted by $\operatorname{id}_{R}(M)$ and $\operatorname{Gid}_{R}(M)$.
+In addition, for an $R$-module $M$, we let $\operatorname{H-dim}_{R}(M)$ stand for $\operatorname{pd}_{R}(M)$ or $\operatorname{G-dim}_{R}(M)$, and $\operatorname{Hid}_{R}(M)$ for the corresponding injective version.
+If $\boldsymbol{X}$ and $\boldsymbol{Y}$ are complexes in $\mathcal{D}(R)$, we set
+$\operatorname{Ext}_{R}^{i}(\boldsymbol{X},\boldsymbol{Y})=\operatorname{H}_{-i}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{X},\boldsymbol{Y}))$ and $\operatorname{Tor}_{i}^{R}(\boldsymbol{X},\boldsymbol{Y})=\operatorname{H}_{i}(\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\boldsymbol{Y}),\text{ for all }i.$
+2. Finite homological dimensions of Ext of complexes
+In this section, we derive results regarding the finiteness of homological dimension of Ext of complexes and conclude the corresponding results for Ext of modules.
+Lemma 2.1.
+Let $\boldsymbol{X}$ be a complex in $\mathcal{D}_{\square}(R).$
   1. (1)
-    $\mathcal{C}_{\sqsubset}(R)$ (resp. $\mathcal{C}_{\sqsupset}(R)$) denotes the category of complexes bounded to the left (resp. right).
+    If $\operatorname{Gid}_{R}(\operatorname{H}_{i}(\boldsymbol{X}))<\infty$ for all $i$, then $\operatorname{Gid}_{R}(\boldsymbol{X})<\infty$.
   2. (2)
-    $\mathcal{C}_{(\sqsubset)}(R)$ (resp. $\mathcal{C}_{(\sqsupset)}(R)$) denotes the category of homologically bounded complexes to the left (resp. to the right).
-  3. (3)
-    $\mathcal{C}_{(\square)}(R)$ denotes the category of homologically bounded complexes.
-  4. (4)
-    $\mathcal{C}^{\rm P}(R)$ (resp. $\mathcal{C}^{\rm I}(R)$) denotes the category of complexes of $R$-modules of projective modules (resp. injective modules).
-  5. (5)
-    $\mathcal{C}^{(\rm f)}(R)$ denote the category of complexes with finitely generated homologies.
-Moreover, for a subcategory $\mathcal{S}(R)\subseteq\mathcal{C}(R)$ we set $\mathcal{S}^{(\rm f)}(R)=\mathcal{S}(R)\cap\mathcal{C}^{(\rm f)}(R)$.
-2.3. Shift
-Let $n$ be an integer. For a complex $\boldsymbol{X},$ its $n$-fold shift $\Sigma^{n}\boldsymbol{X}$ is defined by:
-$(\Sigma^{n}\boldsymbol{X})_{i}=X_{i-n}\,\,\text{ and }\,\,\partial_{i}^{\Sigma^{n}\boldsymbol{X}}=(-1)^{n}\partial_{i-n}^{\boldsymbol{X}}.$
-2.4. Hom and tensor of complexes
-Let $\boldsymbol{X}$ and $\boldsymbol{Y}$ be complexes of $R$-modules. The Hom complex $\operatorname{Hom}_{R}(\boldsymbol{X},\boldsymbol{Y})$ is defined by:
-$\operatorname{Hom}_{R}(\boldsymbol{X},\boldsymbol{Y})_{i}=\prod_{j\in\mathbb{Z}}\operatorname{Hom}_{R}(X_{j},Y_{i+j})$
-and
-$\partial_{i}^{\text{Hom}_{R}(\boldsymbol{X},\boldsymbol{Y})}((\psi_{j}))=(\partial_{j}^{\boldsymbol{Y}}\circ\psi_{j}-(-1)^{i}\psi_{p-1}\circ\partial_{j}^{\boldsymbol{X}})_{j}$
-The tensor product complex $\boldsymbol{X}\otimes_{R}\boldsymbol{Y}$ is defined by:
-$\left(\boldsymbol{X}\otimes_{R}\boldsymbol{Y}\right)_{i}=\bigoplus_{j\in\mathbb{Z}}X_{j}\otimes_{R}Y_{i-j}$
-and
-$\partial_{i}^{\boldsymbol{X}\otimes_{R}\boldsymbol{Y}}\left(m_{j}\otimes n_{i-j}\right)=\partial_{i}^{\boldsymbol{X}}\left(m_{j}\right)\otimes n_{i-j}+(-1)^{j}m_{j}\otimes\partial_{i-j}^{\boldsymbol{Y}}\left(n_{i-j}\right).$
-2.5. Derived category
-We denote by $\mathcal{D}(R)$ the derived category of $R$.
-This is the category of $R$-complexes localized at the class of all quasi-isomorphisms.
-Describing this category can be technical, so for more details, we refer the reader to [10].
-Nevertheless, we mention the following facts:
-  1. (1)
-    The objects of $\mathcal{D}(R)$ are exactly the objects of $\mathcal{C}(R)$.
-  2. (2)
-    A morphism in $\mathcal{C}(R)$ is a morphism in $\mathcal{D}(R)$, and it is an isomorphism in $\mathcal{D}(R)$ if and only if it is a quasi-isomorphism in $\mathcal{C}(R)$.
-By $\simeq$ we indicate isomorphism in $\mathcal{D}(R)$.
-2.6. Derived functors
-The symbols $\operatorname{\mathbf{R}Hom}(-,-)$ and $-\otimes_{R}^{\mathbf{L}}-$ denote, respectively, the right and left derived functors of the homomorphism and tensor product functors for complexes over $R$.
-They are bifunctors on the derived category of $R$, defined up to isomorphism in $\mathcal{D}(R)$.
-The bifunctor $\operatorname{\mathbf{R}Hom}(-,-)$ satisfies the following: Let $\boldsymbol{X}$ and $\boldsymbol{Y}$ be complexes.
-  1. (1)
-    If $\boldsymbol{X}\in\mathcal{C}_{(\sqsupset)}(R)$ and if $\boldsymbol{P}$ is a complex in $\mathcal{C}_{\sqsupset}^{\rm P}(R)$ such that $\boldsymbol{X}\simeq\boldsymbol{P}$, then $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{X},\boldsymbol{Y})=\operatorname{Hom}_{R}(\boldsymbol{P},\boldsymbol{Y})$.
-  2. (2)
-    If $\boldsymbol{Y}\in\mathcal{C}_{(\sqsubset)}(R)$ and if $\boldsymbol{I}$ is a complex in $\operatorname{C}_{\sqsubset}^{\rm I}(R)$ such that $\boldsymbol{I}\simeq\boldsymbol{Y}$, then $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{X},\boldsymbol{Y})=\operatorname{Hom}_{R}(\boldsymbol{X},\boldsymbol{I})$.
-Further, the bifunctor $-\otimes_{R}^{\mathbf{L}}-$ satisfies the following: Let $\boldsymbol{X}$ and $\boldsymbol{Y}$ be complexes.
-  1. (1)
-    If $\boldsymbol{X}\in\mathcal{C}_{(\sqsupset)}(R)$ and if $\boldsymbol{P}$ is a complex in $\mathcal{C}_{\sqsupset}^{\rm P}(R)$ such that $\boldsymbol{X}\simeq\boldsymbol{P}$ , then $\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\boldsymbol{Y}=\boldsymbol{P}\otimes_{R}\boldsymbol{Y}$.
-  2. (2)
-    If $\boldsymbol{Y}\in\mathcal{C}_{(\sqsupset)}(R)$ and if $\boldsymbol{P}$ is a complex in $\mathcal{C}_{\sqsupset}^{\rm P}(R)$ such that $\boldsymbol{Y}\simeq\boldsymbol{P}$, then $\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\boldsymbol{Y}=\boldsymbol{X}\otimes_{R}\boldsymbol{P}$.
-If $\boldsymbol{X}$ and $\boldsymbol{Y}$ are complexes, we set
-$\operatorname{Ext}_{R}^{i}(\boldsymbol{X},\boldsymbol{Y})=\operatorname{H}_{-i}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{X},\boldsymbol{Y}))$ and $\operatorname{Tor}_{i}^{R}(\boldsymbol{X},\boldsymbol{Y})=\operatorname{H}_{i}(\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\boldsymbol{Y}),\text{ for all }i.$
-2.7. Standard isomorphisms
-Let $\boldsymbol{X},\boldsymbol{Y}$ and $\boldsymbol{Z}$ be complexes. Then we have the following isomorphisms in $\mathcal{D}(R)$:
-  1. (Com)
-    $\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\boldsymbol{Y}\simeq\boldsymbol{Y}\otimes_{R}^{\mathbf{L}}\boldsymbol{L}$.
-  2. (Asoc)
-    $(\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\boldsymbol{Y})\otimes_{R}^{\mathbf{L}}\boldsymbol{Z}\simeq\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\left(\boldsymbol{Y}\otimes_{R}^{\mathbf{L}}\boldsymbol{Z}\right).$
-  3. (Adj)
-    $\operatorname{\mathbf{R}Hom}_{R}\left(\boldsymbol{X}\otimes_{R}^{\mathbf{L}}\boldsymbol{Y},\boldsymbol{Z}\right)\simeq\operatorname{\mathbf{R}Hom}_{R}\left(\boldsymbol{X},\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{Y},\boldsymbol{Z})\right).$
-  4. (Swap)
-    $\operatorname{\mathbf{R}Hom}_{R}\left(\boldsymbol{X},\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{Y},\boldsymbol{Z})\right)\simeq\operatorname{\mathbf{R}Hom}_{R}\left(\boldsymbol{Y},\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{X},\boldsymbol{Z})\right).$
-2.8. Homological dimensions
-(1) The projective dimension of a complex $\boldsymbol{X}\in\mathcal{C}_{(\sqsupset)}(R)$ is defined as:
-$\operatorname{pd}_{R}(\boldsymbol{X})=\inf\{\sup\{i\in\mathbb{Z}:P_{i}\not=0\}:\boldsymbol{X}\simeq\boldsymbol{P}\in\mathcal{C}_{\sqsupset}^{\rm P}(R)\}.$
-(2) The injective dimension of a complex $\boldsymbol{Y}\in\mathcal{C}_{(\sqsupset)}(R)$ is defined as:
-$\operatorname{id}_{R}(\boldsymbol{Y})=\inf\{\sup\{i\in\mathbb{Z}:I_{-i}\not=0\}:\boldsymbol{Y}\simeq\boldsymbol{I}\in\mathcal{C}_{\sqsubset}^{\rm I}(R)\}.$
-(3) A finitely generated $R$-module $G$ is totally reflexive if $G$ is reflexive and $\operatorname{Ext}_{R}^{i}(G,R)=\operatorname{Ext}_{R}^{i}(G^{\ast},R)=0$ for all $i>0$.
-By $\mathcal{C}^{\operatorname{G}}(R)$ we denote the category of complexes of totally reflexives $R$-modules.
-The Gorenstein dimension of a complex $\boldsymbol{X}\in\mathcal{C}_{(\sqsupset)}^{({\rm f})}(R)$ is defined as
-$\operatorname{G-dim}_{R}(\boldsymbol{X})=\inf\{\sup\{i\in\mathbb{Z}:G_{i}\not=0\}:\boldsymbol{X}\simeq\boldsymbol{G}\in\mathcal{C}_{\sqsupset}^{\rm G}(R)\}.$
-(4) An $R$-module $M$ is said to be $G$-injective if there exists an exact complex of injective $R$-modules such that
-$\boldsymbol{I}:\cdots\xrightarrow{d_{2}}{I_{2}}\xrightarrow{d_{1}}{I_{0}}\xrightarrow{d_{0}}I_{-1}\xrightarrow{d-1}\cdots$
-such that $M\cong\ker(d_{0})$ and the complex $\operatorname{Hom}_{R}(J,\boldsymbol{I})$ is also exact for any injective $R$-module $J$.
-By $\mathcal{C}^{\operatorname{GI}}(R)$ we denote the category of complexes of Gorenstein injective modules.
-The Gorenstein injective dimension of a complex $\boldsymbol{Y}\in\mathcal{C}_{(\sqsubset)}(R)$ is defined as:
-$\operatorname{Gid}_{R}(\boldsymbol{Y})=\inf\{\sup\{i\in\mathbb{Z}:I_{-i}\not=0\}:\boldsymbol{Y}\simeq\boldsymbol{I}\in\mathcal{C}_{\sqsubset}^{\rm GI}(R)\}.$
-Theorem 2.1.
-The following class of $\mathcal{C}(R)$ are closed under short exact sequences.
-  1. (1)
-    ([10, Corollary 8.1.9]) The class of complexes in $\mathcal{C}_{(\sqsupset)}(R)$ with finite projective dimension.
-  2. (2)
-    ([10, Corollary 8.2.9])  The class of complexes in $\mathcal{C}_{(\sqsubset)}(R)$ with finite injective dimension.
-  3. (3)
-    ([6, Lemma 2.1.12 & Corollary 2.3.8]) The class of complexes in $\mathcal{C}_{(\sqsupset)}^{(\rm f)}(R)$ with finite Gorenstein dimension.
-  4. (4)
-    ([6, Lemma 3.2.12 & Corollary 6.2.5])  The class of complexes in $\mathcal{C}_{(\sqsubset)}(R)$ with finite Gorenstein injective dimension.
-Theorem 2.2.
-Let $R$ be a local ring and $\boldsymbol{X}$ be a complex.
-  1. (1)
-    ([10, Proposition 18.3.9]) If $\boldsymbol{X}\in\mathcal{C}_{(\sqsupset)}^{(\rm f)}(R)$, then $\operatorname{pd}_{R}(\boldsymbol{X})=\operatorname{pd}_{\widehat{R}}(\boldsymbol{X}\otimes_{R}\widehat{R}).$
-  2. (2)
-    ([10, Proposition 18.3.10]) If $\boldsymbol{X}\in\mathcal{C}_{(\sqsubset)}^{(\rm f)}(R)$, then $\operatorname{id}_{R}(\boldsymbol{X})=\operatorname{id}_{\widehat{R}}(\boldsymbol{X}\otimes_{R}\widehat{R}).$
-  3. (3)
-    ([23, 2.4]) If $\boldsymbol{X}\in\mathcal{C}_{(\square)}^{(\rm f)}(R)$, then $\operatorname{G-dim}_{R}(\boldsymbol{X})=\operatorname{G-dim}_{\widehat{R}}(\boldsymbol{X}\otimes_{R}\widehat{R}).$
-  4. (4)
-    ([10, Theorem 19.2.34]) If $\boldsymbol{X}\in\mathcal{C}_{(\square)}^{(\rm f)}(R)$ and $\operatorname{Gid}_{R}(\boldsymbol{X})<\infty$, then $\operatorname{Gid}_{R}(\boldsymbol{X})=\operatorname{Gid}_{\widehat{R}}(\boldsymbol{X}\otimes_{R}\widehat{R}).$
-Proposition 2.3 ([4, Theorem 8.7]).
-Let $R$ be a local ring and let $M$ be a finitely generated $R$-module. Then:
-  1. (1)
-    If $\operatorname{H\text{-}dim}_{R}(M)<\infty$, then $\operatorname{H\text{-}dim}_{R}(M)+0pt(M)=0pt(R)$.
-  2. (2)
-    For every integer $n\geq 0$, $\operatorname{H\text{-}dim}_{R}(\Omega^{n}(M))=\max\{\operatorname{H\text{-}dim}_{R}(M)-n,\,0\}.$
-  3. (3)
-    $\operatorname{G-dim}_{R}(M)\leq\operatorname{pd}_{R}(M)$ with equality whenever $\operatorname{pd}_{R}(M)$ is finite.
-2.9. Semidualizing complexes.
-An $R$-complex $\boldsymbol{C}\in\mathcal{C}_{(\square)}^{(\rm f)}(R)$ is semidualizing if the natural morphism $R\to\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{C},\boldsymbol{C})$ is an isomorphism in $\mathcal{D}(R)$.
-An $R$-module $C$ is said to be semidualizing if it corresponds to a semidualizing complex in $\mathcal{C}_{(\square)}^{(\rm f)}(R)$ ([7]).
-2.10. Auslander categories
-Let $\boldsymbol{C}$ be a semidualizing complex.
-  1. (1)
-    The Auslander class of $R$, denoted by $\mathcal{A}_{\boldsymbol{C}}(R)$, is the full subcategory of $\mathcal{D}(R)$ composed for all the complex $\boldsymbol{X}\in\mathcal{C}_{(\square)}(R)$ such that $\boldsymbol{C}\otimes_{R}^{\mathbf{L}}\boldsymbol{X}\in\mathcal{C}_{(\square)}(R)$ and the canonical map
-    $\boldsymbol{X}\to\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{C},\boldsymbol{C}\otimes_{R}^{\mathbf{L}}\boldsymbol{X})$
-    is an isomorphism in $\mathcal{D}(R)$.
-  2. (2)
-    The Bass class of $R$, denoted by $\mathcal{B}_{\boldsymbol{C}}(R)$, is the full subcategory of $\mathcal{C}_{(\square)}(R)$ composed for all the complexes $\boldsymbol{Y}\in\mathcal{C}_{(\square)}(R)$ such that $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{C},\boldsymbol{Y})\in\mathcal{C}_{(\square)}(R)$ and the canonical map
-    $\boldsymbol{C}\otimes_{R}^{\mathbf{L}}\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{C},\boldsymbol{Y})\to\boldsymbol{Y}$
-    is an isomorphism in $\mathcal{D}(R)$.
-Denote by $\mathcal{F}(R)$ and $\mathcal{I}(R)$, respectively, the subcategories of $\mathcal{D}(R)$ consisting of complexes isomorphic to homologically bounded complexes of flat $R$-modules and of injective $R$-modules.
-Then $\mathcal{F}(R)\subset\mathcal{A}_{\boldsymbol{C}}(R)$ and $\mathcal{I}(R)\subset\mathcal{B}_{\boldsymbol{C}}(R)$.
-2.11. Dualizing complex
-Assume that $R$ is a local ring.
-A dualizing complex $\boldsymbol{D}$ is a bounded complex of injective $R$-modules with finitely generated homologies such that the natural morphism $R\rightarrow\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{D})$ is an isomorphism in $\mathcal{D}(R)$.
-The ring $R$ has a dualizing complex if and only if $R$ is a homomorphic image of a Gorenstein local ring ([6, A.8.3.2] and [24, Corollary 1.4]).
-In particular, if $R$ is complete, then $R$ admits a dualizing complex.
-If $\boldsymbol{D}$ and $\boldsymbol{D}^{\prime}$ are dualizing complexes, then $\boldsymbol{D}\simeq\Sigma^{n}\boldsymbol{D}$ for some integer $n$ ([6, A.8.3.3]).
-If $\boldsymbol{X}$ is a complex in $\mathcal{C}_{(\square)}^{(\rm f)}(R)$, then $\boldsymbol{X}\simeq\operatorname{\mathbf{R}Hom}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{X},\boldsymbol{D}),\boldsymbol{D})$ ([6, A.8.5]).
-Proposition 2.4.
-Assume $R$ is a local ring and let $\boldsymbol{D}$ be a dualizing complex.
-  1. (1)
-    ([6, Corollary 3.1.11]) A complex $\boldsymbol{X}$ in $\mathcal{C}_{(\sqsupset)}^{(\rm f)}(R)$ has finite Gorenstein dimension if and only if $\boldsymbol{X}\in\mathcal{A}_{\boldsymbol{D}}(R)$.
-  2. (2)
-    ([9, Theorem 4.4]) A complex $\boldsymbol{X}$ in $\mathcal{C}_{(\sqsubset)}(R)$ has finite Gorenstein injective dimension if and only if $\boldsymbol{X}\in\mathcal{B}_{\boldsymbol{D}}(R)$.
-A celebrated result that relates to the Auslander categories, and the subcategories $\mathcal{F}(R)$ and $\mathcal{I}(R)$ is the following, which is known as the Foxby Equivalence.
-Theorem 2.5 ([6, Theorem 3.3.2]).
-Assume that $R$ is a local ring with a dualizing complex $\boldsymbol{D}$. Let $\boldsymbol{X}$ and $\boldsymbol{Y}$ be complexes in $\mathcal{C}_{(\square)}(R)$. Then:
-$\boldsymbol{X}\in\mathcal{A}_{\boldsymbol{D}}(R)\quad$ $\Longleftrightarrow\quad\boldsymbol{D}\otimes_{R}^{\mathbf{L}}\boldsymbol{X}\in\mathcal{B}_{\boldsymbol{D}}(R);$
-$\boldsymbol{Y}\in\mathcal{B}_{\boldsymbol{D}}(R)\quad$ $\Longleftrightarrow\quad\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{Y})\in\mathcal{A}_{\boldsymbol{D}}(R);$
-$\boldsymbol{X}\in\mathcal{F}(R)\quad$ $\Longleftrightarrow\quad\boldsymbol{D}\otimes_{R}^{\rm\mathbf{L}}\boldsymbol{X}\in\mathcal{I}(R);$
-$\boldsymbol{Y}\in\mathcal{I}(R)\quad$ $\Longleftrightarrow\quad\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{Y})\in\mathcal{F}(R).$
-In complementation to Theorem 2.2(4), one has the following.
-Theorem 2.6 ([7, Proposition 5.9] & [10, Theorem 19.2.34]).
-Let $R$ be a local ring with a dualizing complex and let $\boldsymbol{X}\in\mathcal{C}_{(\square)}^{(\rm f)}$.
-Then $\operatorname{Gid}_{R}(\boldsymbol{X})=\operatorname{Gid}_{\widehat{R}}(\boldsymbol{X}\otimes_{R}\widehat{R})$.
-2.12. Stability results
-We close this section with the following results, which will be used in the next section.
-Theorem 2.7.
-Assume $R$ is a local ring, and let $\boldsymbol{X}$ and $\boldsymbol{P}$ two complexes in $\mathcal{C}_{(\square)}^{(\rm f)}(R)$.
-Suppose that $\boldsymbol{P}$ is non-acylic and that $\operatorname{pd}_{R}(\boldsymbol{P})<\infty$.
-Then:
-  1. (1)
-    ([23, Theorems 5.6 & 5.8]) $\operatorname{H-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{P},\boldsymbol{X}))<\infty\Longleftrightarrow\operatorname{H-dim}_{R}(\boldsymbol{X})<\infty$.
-  2. (2)
-    ([23, Theorems 5.1 & 5.7]) $\operatorname{H-dim}_{R}(\boldsymbol{P}\otimes_{R}^{\mathbf{L}}\boldsymbol{X})<\infty$Longleftrightarrow $\operatorname{H-dim}_{R}(\boldsymbol{X})<\infty$.
-Theorem 2.8.
-([9, Corollary 6.4]) Assume $R$ is a local ring with a dualizing complex $\boldsymbol{D}$. Then for a complex $\boldsymbol{X}\in\mathcal{C}_{(\square)}^{(\rm f)}(R)$ one has:
-$\operatorname{Gid}_{R}(\boldsymbol{X})<\infty\Longleftrightarrow\operatorname{G-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{X},\boldsymbol{D}))<\infty.$
-Theorem 2.9.
-Assume $R$ is a local ring, and let $\boldsymbol{Y}$ and $\boldsymbol{P}$ be complexes in $\mathcal{C}_{(\square)}^{(\rm f)}(R)$.
-Suppose that $\boldsymbol{P}$ is non-acyclic and that $\operatorname{pd}_{R}(\boldsymbol{P})<\infty$.
-In the case $\operatorname{H-dim}_{R}=\operatorname{G-dim}_{R}$, assume further that $R$ has a dualizing complex.
-  1. (1)
-    If $\operatorname{Hid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{P},\boldsymbol{Y}))<\infty$, then $\operatorname{Hid}_{R}(\boldsymbol{Y})<\infty$.
-  2. (2)
-    If $\operatorname{Hid}_{R}(\boldsymbol{Y}\otimes_{R}^{\mathbf{L}}\boldsymbol{P})<\infty$, then $\operatorname{Hid}_{R}(\boldsymbol{Y})<\infty$.
-Proof.
-If $\operatorname{H-dim}=\operatorname{pd}$, it follows from [8, Lemmas 6.2.10 & 6.2.15].
-On the other hand, if $\operatorname{H-dim}=\operatorname{G-dim}$, then the theorem follows from [32, Proposition 4.1(b)] and Proposition 2.4(2).
-∎
-3. Finite homological dimensions of Ext of complexes
-In this section, we derive results regarding the finiteness of homological dimension of Ext of complexes, and conclude the corresponding results for Ext of modules.
-The motivation arises from the [37, Section 6], where the respective authors obtained a result concerning finiteness of the injective dimension of Ext of complexes ([37, Proposition 4.2]) using [37, Lemma 4.1] and Bass Series, and subsequently derived the corresponding result for Ext of modules.
-Inspired by their approach, we first derive counteparts of the lemma mentioned considering other homological dimensions and then use Poincaré Series or stability results for concluding our results.
-Lemma 3.1.
-Let
-$\boldsymbol{X}:\cdots\longrightarrow X_{l+1}\xrightarrow{\partial_{l+1}}X_{l}\xrightarrow{\partial_{l}}X_{l-1}\rightarrow\cdots$
-be a homologically bounded complex. If $\operatorname{Gid}_{R}(\operatorname{H}_{i}(\boldsymbol{X}))<\infty$ for all $i,$ then $\operatorname{Gid}_{R}(\boldsymbol{X})<\infty$.
-Proof.
-The proof is the same as the proof of [37, Lemma 4.1] but now replacing $\operatorname{Gid}$ by $\operatorname{id}$, considering the complex $\boldsymbol{I}$ of the proof as a resolution of Gorenstein injective modules of $\operatorname{H}_{0}(\boldsymbol{X})$, and using Theorem 2.1(4).
-∎
-Lemma 3.2.
-Let
-$\boldsymbol{X}:\cdots\longrightarrow X_{l+1}\xrightarrow{\partial_{l+1}}X_{l}\xrightarrow{\partial_{l}}X_{l-1}\rightarrow\cdots$
-be a complex in $\mathcal{C}_{(\square)}(R)$.
-Suppose further that $\boldsymbol{X}\in\mathcal{C}^{(\rm{f})}(R)$ in the case $\operatorname{H-dim}_{R}=\operatorname{G-dim}$.
     If $\operatorname{H-dim}_{R}(\operatorname{H}_{i}(\boldsymbol{X}))<\infty$ for all $i$, then $\operatorname{H-dim}_{R}(\boldsymbol{X})<\infty$.
 Proof.
-Let $n$ be the number of non-zero integers $i$ such that $\operatorname{H}_{i}(\boldsymbol{X})\not=0$.
-If $n=0,$ then $\boldsymbol{X}$ is acyclic and hence $\operatorname{H-dim}_{R}(\boldsymbol{X})=-\infty<\infty$.
-Now, we show by induction that the lemma holds when $n\geq 1$.
-If $n=1,$ then there exists an unique integer $i$ such that $\operatorname{H}_{i}(\boldsymbol{X})\not=0$.
-Without loss of generality, we may assume $i=0$.
-Set $Z=\operatorname{Coker}(\partial_{1})$.
-Let
-$\boldsymbol{Y}:\cdots\rightarrow 0\rightarrow Z\xrightarrow{\tilde{\partial}_{0}}X_{-1}\xrightarrow{\partial_{-1}}X_{-2}\rightarrow\cdots,$
-where ~$\tilde{\partial}_{0}$ is the map induced of $\partial_{0}$.
-The natural homomorphism $\boldsymbol{X}\to\boldsymbol{Y}$ is a quasi-isomorphism, so $\boldsymbol{X}\simeq\boldsymbol{Y}$ in $\mathcal{D}(R)$.
-Also, we may note that the natural homomorphism $\operatorname{H}_{0}(\boldsymbol{X})\to\boldsymbol{Y}$, is a quasi-isomorphism, so $\operatorname{H}_{0}(\boldsymbol{X})\simeq\boldsymbol{Y}$ in $\mathcal{D}(R)$.
-Thus, $\operatorname{H}_{0}(\boldsymbol{X})\simeq\boldsymbol{X}$ in $\mathcal{D}(R)$.
-As $\operatorname{H-dim}_{R}(\operatorname{H}_{0}(\boldsymbol{X}))<\infty,$ it follows that $\operatorname{H-dim}_{R}(\boldsymbol{X})<\infty$.
-Suppose $n>1$. We may suppose that $\sup\{i\in\mathbb{Z}:\operatorname{H}_{i}(\boldsymbol{X})\neq 0\}=0$. As above, if $Z=\operatorname{Coker}(\partial_{1})$ and $\boldsymbol{Y}$ is the complex
-$\boldsymbol{Y}:\cdots\rightarrow 0\rightarrow Z\xrightarrow{\tilde{\partial}_{0}}X_{-1}\xrightarrow{\partial_{-1}}X_{-2}\rightarrow\cdots,$
-then $\boldsymbol{X}\simeq\boldsymbol{Y}$ in $\mathcal{D}(R)$. Since $\operatorname{H-dim}_{R}(\operatorname{H}_{0}(\boldsymbol{X}))<\infty$, then there exists a bounded complex
-$\boldsymbol{T}:\cdots\rightarrow T_{i+1}\rightarrow T_{i}\rightarrow\cdots\rightarrow T_{1}\rightarrow T_{0}\rightarrow 0$
-such that $\operatorname{H-dim}_{R}(T_{i})=0$ for all $i\geq 0$, and
-$\operatorname{H}_{i}(\boldsymbol{T})=\begin{cases}\operatorname{H}_{0}(\boldsymbol{X}),&\text{if }i=0\\ 0,&\text{if }i>0.\end{cases}$
-Let $\boldsymbol{C}$ be the complex
-$\boldsymbol{C}:\cdots\rightarrow T_{1}\rightarrow T_{0}\rightarrow Z\xrightarrow{\tilde{\partial}_{0}}X_{-1}\xrightarrow{\partial_{-1}}X_{-2}\rightarrow\cdots,$
-where $T_{0}\to Z$ is the composition of the maps $T_{0}\to\operatorname{H}_{0}(X)$ and $\operatorname{H}_{0}(X)\to Z$.
-Let us consider that the complexes $\boldsymbol{Y}$ and $\boldsymbol{C}$ are indexing such that $Y_{-1}=Z$ and $C_{0}=T_{0}$.
-Then we have naturally an exact sequence
-$0\to\boldsymbol{Y}\to\boldsymbol{C}\to\boldsymbol{T}\to 0$
-Note that $\operatorname{H}_{i}(\boldsymbol{C})=0$ for all $i\geq-1$, and $\operatorname{H}_{i}(\boldsymbol{C})=\operatorname{H}_{i+1}(\boldsymbol{X})$ for $i\leq-2$.
-Hence, the number of non-zero homologies of $\boldsymbol{C}$ is $n-1$.
-Thus, by induction, $\operatorname{H-dim}_{R}(\boldsymbol{C})<\infty$.
-Then from Theorem 2.1(1) or (3), we see that $\operatorname{H-dim}_{R}(\boldsymbol{Y})<\infty$.
-Thus, since $\boldsymbol{X}\simeq\boldsymbol{Y}$, we see that $\operatorname{H-dim}_{R}(\boldsymbol{X})<\infty$.
-∎
+We prove item (1).
+This is done by induction on $s=\sup\boldsymbol{X}-\inf\boldsymbol{X}$.
+If $s=0$, then $\boldsymbol{X}\simeq\Sigma^{j}\operatorname{H}_{j}(\boldsymbol{X})$ for some $j$, and the conclusion follows.
+Assume $s>0$ and let $u=\sup\boldsymbol{X}$.
+Note that there is a natural short exact sequence of complexes $0\to\boldsymbol{Y}\to\boldsymbol{X}\to\tau_{{}_{\subset u-1}}(\boldsymbol{X})\to 0,$ where $\tau_{{}_{\subset u-1}}(\boldsymbol{X})$ denotes the (t1)(t-1) soft left truncation of $\boldsymbol{X}$.
+In addition, note $\boldsymbol{X}\to\tau_{{}_{\subset u-1}}(\boldsymbol{X})$ induces an isomorphism in homologies in degree $\leq u-1$, and that $\operatorname{H}_{i}(\tau_{{}_{\subset u-1}}(\boldsymbol{X}))=0$ for i>u1i>u-1.
+Also, for $i\not=u$, we have that $\operatorname{H}_{i}(\boldsymbol{Y})=0$ and $\operatorname{H}_{u}(\boldsymbol{Y})\cong\operatorname{H}_{u}(\boldsymbol{X})$.
+If $\operatorname{Gid}_{R}(\operatorname{H}_{i}(\boldsymbol{X}))<\infty$ for all $i$, note that $\boldsymbol{Y}$ and $\tau_{{}_{\subset t-1}(\boldsymbol{X})}$ satisfy the induction hypothesis, and therefore they have finite Gorenstein injective dimension.
+Thus, the conclusion follows from the fact that the class of complexes in $\mathcal{D}(R)$ with finite Gorenstein injective dimension has the two out three property [10, Proposition 9.2.15].
+The proof of item (2) is similar using [10, Corollary 8.1.9 and Proposition 9.1.16]. ∎
 Theorem 2.2.
-Assume $R$ is local, and let $\boldsymbol{M}$ and $\boldsymbol{N}$ be non-acyclic complexes in $\mathcal{C}_{(\square)}^{\rm(f)}(R)$ such that $\operatorname{pd}_{R}(\boldsymbol{M})<\infty$.
+Assume $R$ is local, and let $\boldsymbol{M}$ and $\boldsymbol{N}$ be non-acyclic complexes in $\mathcal{D}_{\square}^{\rm f}(R)$ such that $\operatorname{pd}_{R}(\boldsymbol{M})<\infty$.
   1. (1)
@@
 Proof.
-We know that $\operatorname{Ext}_{R}^{i}(\boldsymbol{M},\boldsymbol{N})=\operatorname{H}_{-i}(\mathbf{R}\mathrm{Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))$ for all $i$.
-From [6, A.4.4 and A.5.4], ones sees that $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N})\mathcal{C}_{(\square)}^{\rm(f)}(R)$.
-By Lemma 3.2 (resp.
-Lemma 3.1), we conclude $\operatorname{H-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$ (resp.
+We know that $\operatorname{Ext}_{R}^{i}(\boldsymbol{M},\boldsymbol{N})=\operatorname{H}_{-i}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))$ for all $i$.
+From [6, A.4.4 and A.5.4], one sees that $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N})\in\mathcal{D}_{\square}^{\rm f}(R)$.
+By Lemma 2.1(2) (resp.
+Lemma 2.1(1)), we conclude $\operatorname{H-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$ (resp.
 $\operatorname{Gid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$).
-It follows from Theorem 2.7(1) (resp.
-Theorem 2.9(1)) that $\operatorname{H-dim}_{R}(\boldsymbol{N})<\infty$.
-(resp.
+It follows from [19, Theorems 5.6 & 5.8] (resp.
+[28, Proposition 4.1(b)] and [9, Theorem 4.4]) that $\operatorname{H-dim}_{R}(\boldsymbol{N})<\infty$ (resp.
 $\operatorname{Gid}_{R}(\boldsymbol{N})<\infty$), concluding (1) (resp.
@@
 We may assume that $R$ is complete. Then $R$ has a dualizing complex. By Corollary 2.3(2), then $\operatorname{Gid}_{R}(M)<\infty$. Then the result follows from [6, Theorem 6.3.2]. ∎
-Item (2) of the next theorem provides a positive answer to the Gorenstein version of [18, Question 4.5].
+Item (2) of the next theorem provides a positive answer to the Gorenstein version of [15, Question 4.5].
 Theorem 2.5.
-Assume $R$ is local, and let $\boldsymbol{M}$ and $\boldsymbol{N}$ be non-acyclic complexes in $\mathcal{C}_{(\square)}^{(\rm f)}(R)$.
+Assume $R$ is local, and let $\boldsymbol{M}$ and $\boldsymbol{N}$ be non-acyclic complexes in $\mathcal{D}_{\square}^{\rm f}(R)$.
 Suppose that $\operatorname{id}_{R}(\boldsymbol{N})<\infty$.
@@
 We know that $\operatorname{Ext}_{R}^{i}(\boldsymbol{M},\boldsymbol{N})=\operatorname{H}_{-i}(\mathbf{R}\mathrm{Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))$.
-Additionally, by [6, A.4.4 and A.5.2], we see that $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N})\in\mathcal{C}_{(\square)}^{(\rm f)}(R)$.
-In the case (1) (resp.
-(2)), by assumption, $\operatorname{H-dim}_{R}(\operatorname{H}_{i}(\mathbf{R}\mathrm{Hom}_{R}(\boldsymbol{M},\boldsymbol{N})))<\infty$ (resp.
-$\operatorname{Gid}_{R}(\operatorname{H}_{i}(\mathbf{R}\mathrm{Hom}_{R}(\boldsymbol{M},\boldsymbol{N})))<\infty$) for all $i$.
-Thus, by Lemma 3.1 (resp.
-Lemma 3.2), we have that $\operatorname{H-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$ (resp.
-$\operatorname{Hid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$) for all $i$.
-The conclusion of item (1) for $\operatorname{H-dim}_{R}=\operatorname{pd}_{R}$ it follows from [8, Lemma 6.2.12].
-It is remaining to prove (1) for $\operatorname{H-dim}_{R}=\operatorname{G-dim},$ and (2).
-For proving both cases, either by Theorems 2.2, or by assumption and Theorem 2.6, we can let $\boldsymbol{D}$ be a dualizing complex of $R$.
-Since $\operatorname{id}_{R}(\boldsymbol{N})<\infty$, then $\boldsymbol{N}\in\mathcal{B}_{\boldsymbol{D}}(R)$ and hence $\operatorname{pd}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{N}))<\infty$ by the Foxby equivalence.
+Additionally, by [6, A.4.4 and A.5.2], we see that $\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N})\in\mathcal{D}_{\square}^{\rm f}(R)$.
+In the case (1), for all $i$, $\operatorname{H-dim}_{R}(\operatorname{H}_{i}(\mathbf{R}\mathrm{Hom}_{R}(\boldsymbol{M},\boldsymbol{N})))<\infty$.
+Thus, by Lemma 2.1(1), we have that $\operatorname{H-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$.
+As to (2), one considers Lemma 2.1(2) instead so that $\operatorname{Gid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$.
+The conclusion of item (1) for $\operatorname{H-dim}_{R}=\operatorname{pd}_{R}$ follows from [8, Lemma 6.2.12].
+It remains to prove (1) for $\operatorname{H-dim}_{R}=\operatorname{G-dim},$ and (2).
+For proving both, either by [19, 2.4], or by assumption, [7, Proposition 5.9] and [10, Theorem 19.2.34], we can let $\boldsymbol{D}$ be a dualizing complex of $R$.
+Since $\operatorname{id}_{R}(\boldsymbol{N})<\infty$, then $\boldsymbol{N}\in\mathcal{B}(R)$ and hence $\operatorname{pd}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{N}))<\infty$ by the Foxby equivalence.
 We have
-$\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N})$ $\simeq\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{D}\otimes_{R}^{\mathbf{L}}\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{N}))$ because $\text{because }\boldsymbol{N}\in\mathcal{B}_{\boldsymbol{D}}(R)$
+$\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N})$ $\simeq\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{D}\otimes_{R}^{\mathbf{L}}\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{D},\boldsymbol{N}))$ $\text{because }\boldsymbol{N}\in\mathcal{B}(R)$
 […]
-In the case (1) for $\operatorname{H-dim}_{R}=\operatorname{G-dim}_{R}$, we have that $\operatorname{G-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$, concluding from Theorem 2.7(2) that $\operatorname{G-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{D}))<\infty$.
-Hence $\operatorname{Gid}_{R}(\boldsymbol{M})<\infty$ by Theorem 2.8.
+In the case (1) for $\operatorname{H-dim}_{R}=\operatorname{G-dim}_{R}$, we have that $\operatorname{G-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$, concluding from [19, Theorems 5.1] that $\operatorname{G-dim}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{D}))<\infty$.
+Hence $\operatorname{Gid}_{R}(\boldsymbol{M})<\infty$ by [9, Corollary 6.4].
 As to (2), we have that $\operatorname{Gid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{N}))<\infty$.
-Then from the last isomorphism and Theorem 2.9(2), we conclude that $\operatorname{Gid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{D}))<\infty$.
-Since $M\simeq\operatorname{\mathbf{R}Hom}_{R}(\operatorname{\mathbf{R}Hom}_{R}(M,\boldsymbol{D}),\boldsymbol{D})$ by [6, A.8.5], then it follows from Theorem 2.8 that $\operatorname{G-dim}_{R}(\boldsymbol{M})<\infty$.
+It follows from the last isomorphism, [28, Proposition 4.1(b)] and [9, Theorem 4.4] that $\operatorname{Gid}_{R}(\operatorname{\mathbf{R}Hom}_{R}(\boldsymbol{M},\boldsymbol{D}))<\infty$.
+Since $M\simeq\operatorname{\mathbf{R}Hom}_{R}(\operatorname{\mathbf{R}Hom}_{R}(M,\boldsymbol{D}),\boldsymbol{D})$ by [6, A.8.5], then it follows from [9, Corollary 6.4] that $\operatorname{G-dim}_{R}(\boldsymbol{M})<\infty$.
+∎
 Corollary 2.6.
@@
     If $\operatorname{Gid}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t-\operatorname{depth}(M)$, then $\operatorname{G-dim}_{R}(M)<$$\infty$.
-When $\operatorname{H-dim}_{R}=\operatorname{pd}_{R}$, note that Corollary 3.7(1) recovers the celebrated theorem given by Foxby in [15, Corollary 4.4], which states that if there exists a (finitely generated) non-zero $R$-module $M$ with both finite injective dimension and projective dimension, then $R$ is Gorenstein.
-On the other hand, via Corollary 3.7(2), we obtain the following Gorenstein criterion that generalizes [37, Corollary 4.4] and [18, Corollary 4.7].
+When $\operatorname{H-dim}_{R}=\operatorname{pd}_{R}$, note that Corollary 2.6(1) recovers the celebrated theorem of Foxby in [14, Corollary 4.4], which states that if there exists a (finitely generated) non-zero $R$-module $M$ with both finite injective dimension and projective dimension, then $R$ is Gorenstein.
+On the other hand, via Corollary 2.6(2), we obtain the following Gorenstein criterion that generalizes [15, Corollary 4.7] and [32, Corollary 4.4].
 Corollary 2.7.
@@
 Setup.
-From now until the end of the text, $R$ is local, and unless otherwise stated, all $R$-modules will be finitely generated.
-In this section, we derive some results about the implication of having modules of deficiency with certain finite homological dimensions.
-For this, we assume here that $R$ has a dualizing complex $\boldsymbol{D}$.
-Moreover, we suppose that $\boldsymbol{D}$ is normalized in the sense that $\operatorname{Hom}_{R}(k,\boldsymbol{D})\simeq\Sigma^{d}k$, where $d$ is the dimension of $R$.
-The existence of $\boldsymbol{D}$ is equivalent to $R$ being a factor of a Gorenstein local ring $S$, i.e., $R$ is an epimorphic image of $S$.
+In the rest of the paper, $R$ is local, and unless otherwise stated, all $R$-modules will be finitely generated.
+Moreover, in this section, we assume that $R$ has a dualizing complex $\boldsymbol{D}$.
+Also, we assume that $\boldsymbol{D}$ is normalized in the sense that $\sup D=\dim R$.
+The existence of $\boldsymbol{D}$ is equivalent to $R$ being a factor of a Gorenstein local ring $S$, i.e., $R$ is a homomorphic image of $S$.
 In that context, Schenzel [29] generalized the notion of a canonical module of a ring in the following sense: For an $R$-module $M$, set
@@
 The $R$-module $K(M):=K^{\dim_{R}(M)}(M)$ is called the canonical module of $M$.
-The modules of deficiency satisfy the following isomorphisms:
-$H^{i}_{\mathfrak{m}}(M)\cong\text{Hom}_{R}(K^{i}(M),E_{R})\text{ and }K^{i}(M)\cong\operatorname{H}_{i-d}(\operatorname{\mathbf{R}Hom}_{R}(M,\boldsymbol{D})),$
-where $E_{R}$ is the hull envelope of the residue field of $R$.
-Note that the first isomorphism shows that $K^{i}(M)=0$ for $i<0pt(M)$ or $i>\dim(M)$, and $K^{i}(M)\neq 0$ for $i=0pt(M)$ and $i=\dim(M)$.
-For further details about deficiency modules, we refer the reader to [33, 34].
-In [13, 17], the authors investigated the consequences of the finiteness of classical homological dimensions, namely, projective and injective dimensions, of the deficiency modules and the canonical module.
-Motivated by their work, we apply Theorem 3.6 to extend those results to a broader class of homological dimensions considered in this paper.
+By local duality we have $K^{i}(M)=0$ for $i<0pt(M)$ or $i>\dim(M)$, and $K^{i}(M)\neq 0$ for $i=0pt(M)$ and $i=\dim(M)$.
+The modules of deficiency also satisfy the following isomorphism for all $i$:
+$K^{i}(M)\cong\operatorname{H}_{i}(\operatorname{\mathbf{R}Hom}_{R}(M,\boldsymbol{D})).$
+For further details about deficiency modules, we refer the reader to [29].
+In [12], the authors investigated the consequences of the finiteness of classical homological dimensions, namely, projective and injective dimensions, of the deficiency modules.
+Motivated by their work, we apply Theorem 2.5 to extend those results to Gorenstein dimensions.
 We also emphasize that our approach differs from theirs in both method and perspective.
@@
 Proof.
-For each $i$, we note that $K^{i}(M)=\operatorname{H}_{i-d}(\operatorname{\mathbf{R}Hom}_{R}(M,\boldsymbol{D}))=\operatorname{Ext}_{R}^{d-i}(M,\boldsymbol{D})$.
+For each $i$, we note that $K^{i}(M)=\operatorname{H}_{i}(\operatorname{\mathbf{R}Hom}_{R}(M,\boldsymbol{D}))=\operatorname{Ext}_{R}^{-i}(M,\boldsymbol{D})$.
 Since $K^{i}(M)=0$ for $i<0pt(M)$ or $i>\operatorname{dim}_{R}(M)$, by assumption we see have that all modules $\operatorname{Ext}_{R}^{i}(M,\boldsymbol{D})$ have finite $\operatorname{H-dim}_{R}$ or all modules have finite $\operatorname{Hid}$.
-Then the result follows from [37, Proposition 4.2] and Theorem 3.6.
+Then the result follows from [32, Proposition 4.2] and Theorem 2.5.
 ∎
-The following corollary contains the counterparts of [14, Proposition 3.5(b)] with Gorenstein dimensions.
+The following corollary contains the counterparts of [13, Proposition 3.5(b)] to Gorenstein dimensions.
 Corollary 3.2.
@@
 Proof.
-In view of [33, Theorem 1.14], it is enough to prove (1).
+In view of [29, Theorem 1.14], it is enough to prove (1).
 One implication follows from Theorem 3.1(1).
 For the other, suppose that $\operatorname{Hid}_{R}(M)<\infty$.
-Since $M\cong K(K(M))$, then $K(M)$ is Cohen-Macaulay ([33, Theorem 1.14] again) and from Theorem 4.1(2) it follows that $\operatorname{H-dim}_{R}(K(M))<\infty$.
+Since $M\cong K(K(M))$, then $K(M)$ is Cohen-Macaulay ([29, Theorem 1.14] again) and from Theorem 3.1(2) it follows that $\operatorname{H-dim}_{R}(K(M))<\infty$.
 ∎
-5. On Finite Homological dimension of Ext modules over CM rings
+Corollary 3.3.
+Let $R$ be a Cohen-Macaulay local ring with canonical module $\omega$.
+Assume that $\operatorname{Ext}_{R}^{i}(\omega,R)=0$ for all $1\leq i\leq t$.
+Let $^{\dagger}$ stand for $\operatorname{Hom}_{R}(\_,\omega)$.
+If $\operatorname{Gid}_{R}((\omega^{*})^{\dagger})<\infty$, then $R$ is Gorenstein.
+Proof.
+Note that from the vanishing hypothesis and [18, Theorem 2.1(i)] we see that $\omega^{*}$ is maximal Cohen-Macaulay.
+Since $\operatorname{Gid}_{R}((\omega^{*})^{\dagger})<\infty$, then by Corollary 3.2, $\operatorname{G-dim}_{R}(\omega^{\ast})<\infty$.
+Therefore, the conclusion follows from [24, Corollary 5.13].
+∎
+4. A duality for Ext modules with finite homological dimension
 In this section, we establish a duality of certain Ext modules. The following theorem will be often used in this section.
-Proposition 5.1 (Proposition 2.4 & [6, Theorem 3.4.11]).
+Proposition 4.1 ([9, Theorem 4.4], [6, Corollary 3.1.11 & Theorem 3.4.11] ).
 Assume $R$ is Cohen-Macaulay with canonical module $\omega$, and let $M$ be an $R$-module.
   1. (1)
-    $\operatorname{G-dim}_{R}(M)<\infty$ if and only if $M\in\mathcal{A}_{\omega}(R)$.
+    $\operatorname{G-dim}_{R}(M)<\infty$ if and only if $M\in\mathcal{A}(R)$.
   2. (2)
-    $\operatorname{Gid}_{R}(M)<\infty$ if and only if $M\in\mathcal{B}_{\omega}(R)$.
+    $\operatorname{Gid}_{R}(M)<\infty$ if and only if $M\in\mathcal{B}(R)$.
   3. (3)
@@
     $\operatorname{H-dim}_{R}(M)<\infty$ if and only if $\operatorname{Hid}_{R}(M\otimes_{R}\omega)<\infty$.
-In [12, Corollary 42(3)], although the authors provide an independent proof of Proposition 5.1, this proposition assures that their vanishing hypotheses are not necessary.
+In [11, Corollary 42(3)], although the authors provide this without using Proposition 4.1, this proposition assures that their vanishing assumptions are not necessary.
 Corollary 4.2.
@@
 In particular, $\operatorname{G-dim}(\operatorname{Hom}_{R}(M,N))<\infty$.
-Note that $\operatorname{Hom}_{R}(M,N)$ and $N$ belong to $\mathcal{A}_{\omega}(R)$.
-Then by [27, Lemma 5.1], the map $\Phi_{M,N}$ is an isomorphism.
+Note that $\operatorname{Hom}_{R}(M,N)$ and $N$ belong to $\mathcal{A}(R)$.
+Then by [23, Lemma 5.1], the map $\Phi_{M,N}$ is an isomorphism.
 Hence, it follows from Proposition 4.1(3) that $\operatorname{Hid}_{R}(\operatorname{Hom}_{R}(M,N\otimes_{R}\omega))<\infty$.
@@
 $\cong\operatorname{Hom}_{R}(M,N),$
-where the last isomorphism is because $N\in\mathcal{A}_{\omega}(R).$ Thus, $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(M,N))<\infty$. ∎
-Let $\mathcal{A}$ be an abelian category and let $\mathcal{X}$ a class of objects in $\mathcal{A}$.
-We say that $\mathcal{X}$ is closed under short exact sequences if, in each short exact sequence, the fact that two objects are in $\mathcal{X}$ implies that the third so is.
-We say that $\mathcal{X}$ is closed under isomorphisms if every object isomorphic to $\mathcal{X}$, is in $\mathcal{X}$.
+where the last isomorphism follows since $N\in\mathcal{A}(R).$ Thus, $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(M,N))<\infty$. ∎
+Let $\mathcal{A}$ be an abelian category and let $\mathcal{X}$ be a class of objects in $\mathcal{A}$.
+We say that $\mathcal{X}$ is closed under short exact sequences if the two out of three property holds for short exact sequences of objects in $\mathcal{X}$.
+We say that $\mathcal{X}$ is closed under isomorphisms if every object isomorphic to an object in $\mathcal{X}$ is in $\mathcal{X}$.
 Lemma 4.4.
@@
 For $j=0$, this holds by hypothesis.
-Now suppose $0<j\leq n$ and that $\operatorname{Hom}_{R}(\Omega^{j-1}(M),N))<\infty$.
-We have an exact sequence
-[…]
-which yields
-[…]
-since $\operatorname{Ext}^{1}_{R}(\Omega^{j-1}(M),N)\simeq\operatorname{Ext}^{j}_{R}(M,N)$. Such an exact sequence splits into
-[…]
-and
-[…]
-Since $\mathcal{X}$ is closed under short exact sequences and isomorphisms, it follows that $\operatorname{Hom}_{R}(\Omega^{j}(M),N))\in\mathcal{X}$. ∎
-The following theorem improves [27, Lemma 5.1].
+Now suppose $0<j\leq n$ and that $\operatorname{Hom}_{R}(\Omega^{j-1}(M),N))\in\mathcal{X}$.
+We have an exact sequence ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Omega^{j}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R^{n_{j}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Omega^{j-1}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0,}$ which yields
+${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Hom}_{R}(\Omega^{j-1}(M),N)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N^{n_{j}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Hom}_{R}(\Omega^{j}(M),N)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Ext}^{j}_{R}(M,N)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
+since $\operatorname{Ext}^{1}_{R}(\Omega^{j-1}(M),N)\simeq\operatorname{Ext}^{j}_{R}(M,N)$.
+The result follows by breaking this exact sequence into two short ones and using the fact that $\mathcal{X}$ is closed under short exact sequences and isomorphisms.
+The following theorem improves [23, Lemma 5.1]. For nomenclature on semidualizing modules and their corresponding Auslander classes, we refer to [27].
 Theorem 4.5.
@@
 Proof.
-For $i=0$, the isomorphism holds from [27, Lemma 5.1]. For $i>0$, consider the short exact sequence
-[…]
+For $i=0$, the isomorphism holds from [23, Lemma 5.1].
+For $i>0$, consider the short exact sequence ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Omega^{i}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R^{r_{i}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Omega^{i-1}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
 Such an exact sequence yields the following commutative diagram
-[…]
+${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Hom}_{R}(\Omega^{i-1}(M),N)\otimes C\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{\Omega^{i-1}(M),N}}$${\operatorname{Hom}_{R}(R^{r_{i}},N)\otimes C\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{R^{r_{i}},N}}$${\operatorname{Hom}_{R}(\Omega^{i}(M),N)\otimes C\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{\Omega^{i}(M),N}}$${\operatorname{Ext}^{1}_{R}(\Omega^{i-1}(M),N)\otimes C\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Hom}(\Omega^{i-1}(M),N\otimes_{R}C)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Hom}(R^{r_{i}},N\otimes_{R}C)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Hom}(\Omega^{i}(M),N\otimes_{R}C)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Ext}^{1}_{R}(\Omega^{i-1}(M),N\otimes_{R}C)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
 where the second row is exact.
-As to the first one, since $\operatorname{Ext}^{1}_{R}(\Omega^{i-1}(M),N)\cong\operatorname{Ext}^{i}_{R}(M,N)$, by assumption, [35, 1.9(a)] and Lemma 5.4, all $R$-modules of this exact sequence belong to $\mathcal{A}_{C}(R)$.
-Therefore, the first row turns out to be exact as well by [31, Lemma 3.1.12].
-Additionally, by [27, Lemma 5.1], the vertical maps are isomorphisms.
-Then, from the five lemma, we assure the existence of a natural isomorphism $\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N)\otimes_{R}C\cong\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N\otimes_{R}C)$, from which we conclude the result.
-∎
+As to the first one, since $\operatorname{Ext}^{1}_{R}(\Omega^{i-1}(M),N)\cong\operatorname{Ext}^{i}_{R}(M,N)$, by assumption, [30, 1.9(a)] and Lemma 4.4, all $R$-modules of this exact sequence belong to $\mathcal{A}_{C}(R)$.
+Therefore, the first row turns out to be exact as well by [27, Lemma 3.1.12].
+Additionally, by [23, Lemma 5.1], the vertical maps are isomorphisms. The five lemma settles the desired isomorphism. ∎
+Recall that the classes of modules of finite projective dimension, finite injective dimension, finite Gorenstein dimension, and finite Gorenstein injective dimension are closed under short exact sequences; see [10, Corollaries 8.1.9 and 8.2.9, Proposition 9.2.15] and [6, Corollary 1.2.9].
+With these facts in mind, together with the results established earlier in this section, we are now able to prove our main theorem.
 Theorem 4.6.
-Assume $R$ Cohen-Macaulay with canonical module $\omega$, and let $n$ be a non-negative integer.
+Assume $R$ is Cohen-Macaulay with canonical module $\omega$, and let $n$ be a non-negative integer.
 Let $M$ and $N$ be $R$-modules such that $\operatorname{H-dim}_{R}(N)<\infty$.
-Then the following conditions are equivalents:
+Then the following conditions are equivalent:
   1. (1)
@@
 Proof.
-We first prove the equivalence.
-Suppose that $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ (resp.
-$\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega))<\infty$) for all $0\leq i\leq n$.
-Given $0\leq i\leq n$, we must show that $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega))<\infty$ (resp.
-$\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$).
+Once the equivalence is proved, all the isomorphisms follow from Theorem 4.5, Proposition 4.1, and item (1).
+Now, we prove $(1)\Rightarrow(2)$.
+Suppose that $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq n$.
+Given $0\leq i\leq n$, we must show that $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega))<\infty$.
 The case $i=0$ follows from Proposition 4.3.
 Now, suppose $i>0$.
-Consider the exact sequence
-$0\stackrel{{\scriptstyle~{}}}{{\longrightarrow}}\Omega^{i}(M)\stackrel{{\scriptstyle~{}}}{{\longrightarrow}}R^{r_{i}}\stackrel{{\scriptstyle~{}}}{{\longrightarrow}}\Omega^{i-1}(M)\longrightarrow 0$
-Then it induces an exact sequence
-(5.1) $\scriptstyle{0\longrightarrow\operatorname{Hom}_{R}(\Omega^{i-1}(M),N\otimes_{R}\omega)\longrightarrow\operatorname{Hom}_{R}(R^{r_{i}},N\otimes_{R}\omega)\longrightarrow\operatorname{Hom}_{R}(\Omega^{i}(M),N\otimes_{R}\omega)\longrightarrow\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N\otimes_{R}\omega)\longrightarrow 0}.$
-(resp.
-(5.2) $\scriptstyle{0\longrightarrow\operatorname{Hom}_{R}(\Omega^{i-1}(M),N)\longrightarrow\operatorname{Hom}_{R}(R^{r_{i}},N)\longrightarrow\operatorname{Hom}_{R}(\Omega^{i}(M),N)\longrightarrow\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N)\longrightarrow 0}\mbox{)}.$
+Consider the exact sequence $0\stackrel{{\scriptstyle~}}{{\longrightarrow}}\Omega^{i}(M)\stackrel{{\scriptstyle~}}{{\longrightarrow}}R^{r_{i}}\stackrel{{\scriptstyle~}}{{\longrightarrow}}\Omega^{i-1}(M)\longrightarrow 0.$ Then it induces an exact sequence
+$\scriptscriptstyle{0\longrightarrow\operatorname{Hom}_{R}(\Omega^{i-1}(M),N\otimes_{R}\omega)\longrightarrow\operatorname{Hom}_{R}(R^{r_{i}},N\otimes_{R}\omega)\longrightarrow\operatorname{Hom}_{R}(\Omega^{i}(M),N\otimes_{R}\omega)\longrightarrow\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N\otimes_{R}\omega)\longrightarrow 0}.$
 By Proposition 4.1(4), $\operatorname{Hid}_{R}(N\otimes_{R}\omega)<\infty$.
-Additionally, by Lemma 5.4 and Theorem 2.1, $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(\Omega^{j}(M),N))<\infty$ (resp.
-$\operatorname{Hid}_{R}(\operatorname{Hom}_{R}(\Omega^{j}(M),N\otimes_{R}\omega))<\infty$) for $j=i-1,i$, so by Proposition 5.3, $\operatorname{Hid}_{R}(\operatorname{Hom}_{R}(\Omega^{j}(M),N\otimes_{R}\omega))<\infty$ (resp.
-$\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(\Omega^{j}(M),N))<\infty$) for $j=i-1,i$.
-Thus, the first three terms of the exact sequence (5.1) (resp.
-(5.2)) have finite $\operatorname{Hid},$ (resp.
-$\operatorname{H-dim}_{R}$) and therefore $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N\otimes_{R}\omega))<\infty$ (resp.
-$\operatorname{H-dim}_{R}\operatorname{Ext}^{1}_{R}(\Omega^{i-1}(M),N)<\infty$).
+Additionally, by Lemma 4.4, $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(\Omega^{j}(M),N))<\infty$ for j=i1,ij=i-1,i, so by Proposition 4.3, $\operatorname{Hid}_{R}(\operatorname{Hom}_{R}(\Omega^{j}(M),N\otimes_{R}\omega))<\infty$ for j=i1,ij=i-1,i.
+Thus, the first three terms of the exact sequence above have finite $\operatorname{Hid},$ and therefore $\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N\otimes_{R}\omega))<\infty$.
 The result follows from the fact that $\operatorname{Ext}_{R}^{i}(M,N\otimes_{R}\omega)\cong\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N\otimes_{R}\omega)$.
-Now, we prove the last statement.
-Suppose that (1) holds.
-Then the $R$-modules $N$ and $\operatorname{Ext}_{R}^{i}(M,N)$, for all $i=0,\ldots,n$, have finite Gorenstein dimension.
-Thus, these modules belong to $\mathcal{A}_{\omega}(R)$.
-Hence, the isomorphism required follows from Theorem 5.5.
-∎
+As to $(2)\Rightarrow(1)$, we follow the same lines as before by considering the exact sequence
+$\scriptstyle{0\longrightarrow\operatorname{Hom}_{R}(\Omega^{i-1}(M),N)\longrightarrow\operatorname{Hom}_{R}(R^{r_{i}},N)\longrightarrow\operatorname{Hom}_{R}(\Omega^{i}(M),N)\longrightarrow\operatorname{Ext}_{R}^{1}(\Omega^{i-1}(M),N)\longrightarrow 0}$
+for $0<i\leq n$. ∎
+[…]
 Corollary 4.7.
-Assume $R$ is Cohen-Macaulay, and let $M$ and $N$ be non-zero $R$-modules such that $\operatorname{pd}_{R}(N)<\infty$.
+Assume $R$ is Cohen-Macaulay local, and let $M$ and $N$ be non-zero $R$-modules such that $\operatorname{pd}_{R}(N)<\infty$.
 If $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t-\operatorname{depth}(M)$, then $\operatorname{H-dim}_{R}(M)<\infty$.
+In addition, if $\operatorname{Ext}_{R}^{i}(M,N)=0$ for all $1\leq i\leq t-0pt(M)$, then $\operatorname{H-dim}_{R}(M)=0$.
 Proof.
@@
 Then it follows from Corollary 2.6 that $\operatorname{H-dim}_{R}(M)<\infty$.
-∎
-The following corollary contains a generalization of [12, Corollary 30] and [28, Theorem 4.2$(3)\Rightarrow(2)$].
-Corollary 5.8.
-Assume $R$ is Cohen-Macaulay, and let $M$ and $N$ be non-zero $R$-modules such that $\operatorname{pd}_{R}(N)<\infty$ and $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(M,N))<\infty$.
-If $\operatorname{Ext}_{R}^{i}(M,N)=0$ for all $1\leq i\leq t-0pt(M)$, then $\operatorname{H-dim}_{R}(M)=0$.
-Proof.
-By Corollary 5.7, we have that $\operatorname{H-dim}_{R}(M)<\infty$.
-As $\operatorname{pd}_{R}(N)<\infty$, we have that $\operatorname{Ext}_{R}^{i}(M,N)=0$ for all $i>t-0ptM$ by [1, Theorem 4.13], and that $\operatorname{H-dim}_{R}(M)=\sup\{i\geq 0:\operatorname{Ext}_{R}^{i}(M,N)\not=0\}$ by [30, Theorem 2.6(ii)].
+Now, by the vanishing assumption, as $\operatorname{pd}_{R}(N)<\infty$, we have that $\operatorname{Ext}_{R}^{i}(M,N)=0$ for all i>t0pt(M)i>t-0pt(M) by [1, Theorem 4.13], and that $\operatorname{H-dim}_{R}(M)=\sup\{i\geq 0:\operatorname{Ext}_{R}^{i}(M,N)\not=0\}$ by [26, Theorem 2.6(ii)].
 Therefore, $\operatorname{H-dim}_{R}(M)=0$.
 ∎
-Combining Proposition 3.4 with Corollary 5.7, we have the following byproduct.
+Combining Corollary 2.3 with Corollary 4.7, we have the following byproduct.
 Corollary 4.8.
@@
 Then $\operatorname{pd}_{R}(M)<\infty$ if and only if $\operatorname{pd}_{R}(N)<\infty$.
-Next, we explore the consequences of Theorem 5.6 when in such theorem we consider $N=R$.
+Next, we explore the consequences of Theorem 4.6 when one considers $N=R$.
 Recall that the type of an $R$-module $M$ is defined to be the dimension of the $k$-vector space $\operatorname{Ext}^{r}_{R}(k,M)$, where $r=0pt(M)$, and denoted by $\operatorname{r}(M)$.
@@
 Since such isomorphisms are induced by the natural map $\Phi_{M,N}$ in Proposition 4.3, it follows from [1, Theorem 2.8(b)] that $\operatorname{Tor}^{R}_{1}(\operatorname{Tr}_{R}(\Omega^{i}(M)),\omega)=0$ for all $0\leq i\leq n$.
-Now, by taking the tensor product by the residue field of $R$ and dimensions, the required equality holds from [5, Proposition 3.3.11(c)]. ∎
+Now, by taking the tensor product by the residue field of $R$ and counting dimensions, the required equality holds from [5, Proposition 3.3.11(c)]. ∎
 Corollary 4.10.
 Assume $R$ is Cohen-Macaulay, and let $M$ be an $R$-module of depth $r$ and dimension $s$.
-Suppose that $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $t-s\leq i\leq t-r$ Then:
-  1. (1)
-    $\operatorname{H-dim}_{R}(M)<\infty$.
-  2. (2)
-    $\operatorname{dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))\leq t-i$ for all $i\neq t-s$ and $\dim_{R}(\operatorname{Ext}^{t-s}_{R}(M,R))=s$.
-  3. (3)
-    $\operatorname{r}(M)=\mu(\operatorname{Ext}^{t-r}_{R}(M,R))\operatorname{r}(R).$
-  4. (4)
-    If $M$ has type one, then $R$ is Gorenstein.
+Suppose that $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $t-s\leq i\leq t-r$.
+Then $\operatorname{H-dim}_{R}(M)<\infty$.
+Moreover, $\operatorname{dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))\leq t-i$ for all $i\neq t-s$ and $\dim_{R}(\operatorname{Ext}^{t-s}_{R}(M,R))=s$.
 Proof.
 We may assume that $R$ is complete and therefore $R$ admits a canonical module $\omega$.
-By [5, Corollary 2.1.4], $\operatorname{grade}(M)=t-s$, so $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $i<t-s$ and $\operatorname{Ext}_{R}^{t-s}(M,R)\not=0$.
-Then by Corollary 5.7, $\operatorname{H-dim}_{R}(M)<\infty$ and (1) it is proved.
-On the other hand, since $\operatorname{Ext}_{R}^{i}(M,R)=0$ for $i>t-r$, one can see by assumption that $\operatorname{G-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $i\geq 0$.
-Thus, given $i\geq 0$, by Corollary 5.10,
-$\operatorname{Ext}_{R}^{i}(M,R)\otimes_{R}\omega\cong\operatorname{Ext}_{R}^{i}(M,\omega).$
-It follows that the modules $\operatorname{Ext}^{i}_{R}(M,R)$ and $\operatorname{Ext}^{i}_{R}(M,\omega)$ have the same support and so $\dim_{R}(\operatorname{Ext}_{R}^{i}(M,R))=\dim_{R}(\operatorname{Ext}_{R}^{i}(M,\omega))$.
-In this way, item (2) follows from [33, Lemma 1.9 a) and c)].
-Item (3) follows from Corollary 5.10 and [13, Theorem 3.1], and item (4) follows from (3) and [5, Theorem 3.2.10(b)]. ∎
-Corollary 5.12.
-Assume $R$ is Cohen-Macaulay, and let $M$ be an $R$-module of grade $g$.
-If $M$ is Cohen-Macaulay and $\operatorname{H-dim}(\operatorname{Ext}_{R}^{g}(M,R))<\infty,$ then $M$ is $\operatorname{H}$-perfect.
-Proof.
-By [5, Corollary 2.1.4], $g=t-s$ with $s=\dim_{R}(M)$, so by assumption $\operatorname{G-dim}_{R}(\operatorname{Ext}_{R}^{t-s}(M,R))<\infty$.
-Since $M$ is Cohen-Macaulay, it follows from Corollary 5.11(1), that $\operatorname{H-dim}_{R}(M)<\infty$.
-It follows from the formula Proposition 2.3(2) that $\operatorname{H-dim}_{R}(M)=t-s=g$, i.e., $M$ is $\operatorname{H}$-perfect.
+By [5, Corollary 2.1.4], $\operatorname{grade}(M)=t-s$, so $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all i<tsi<t-s and $\operatorname{Ext}_{R}^{t-s}(M,R)\not=0$.
+Then by Corollary 4.7, $\operatorname{H-dim}_{R}(M)<\infty$.
+On the other hand, since $\operatorname{Ext}_{R}^{i}(M,R)=0$ for i>tri>t-r, one can see by assumption that $\operatorname{G-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $i\geq 0$.
+Thus, given $i\geq 0$, by Corollary 4.9, $\operatorname{Ext}_{R}^{i}(M,R)\otimes_{R}\omega\cong\operatorname{Ext}_{R}^{i}(M,\omega).$ It follows that the modules $\operatorname{Ext}^{i}_{R}(M,R)$ and $\operatorname{Ext}^{i}_{R}(M,\omega)$ have the same support and so $\dim_{R}(\operatorname{Ext}_{R}^{i}(M,R))=\dim_{R}(\operatorname{Ext}_{R}^{i}(M,\omega))$.
+In this way, the last assertion follows from [29, Lemma 1.9 a) and c)].
 ∎
 Remark 4.11.
-In [16, Theorem 2 and 2’], Foxby showed that if $R$ is Cohen-Macaulay, then an $R$-module $M$ of grade $g$ is $\operatorname{H}$-perfect if and only if $M$ is Cohen-Macaulay and $\operatorname{Ext}_{R}^{g}(M,R)$ is $\operatorname{H}$-perfect.
-Note that Corollary 5.12 generalizes the converse of that result.
-In view of [18, Theorem 3.3], it should be noted that Corollary 5.11(3) generalizes the Ghosh-Puthenpurakal type formula [18, Theorem 3.5] for Cohen-Macaulay rings and answers negatively [18, Question 3.9] when we are over Gorenstein rings.
-Corollary 5.14.
-Assume $R$ is Gorenstein. For any $R$-module $M$ one has
-$\operatorname{r}(M)=\mu(\operatorname{Ext}_{R}^{\operatorname{G-dim}_{R}(M)}(M,R)).$
-Although most of the next result is known, our results (including Theorem 5.6) show that such different characterizations of the Gorenstein property of a Cohen-Macaulay ring come from the same place.
-Corollary 5.15.
-Assume $R$ is Cohen-Macaulay and let $\omega$ be its canonical module. We let $\_^{\dagger}$ stands for $\operatorname{Hom}_{R}(\_,\omega)$. The following are equivalent:
-  1. (1)
-    $R$ is Gorenstein.
-  2. (2)
-    $\operatorname{Gid}_{R}(k)<\infty$.
-  3. (3)
-    $\operatorname{Gid}(R)<\infty$.
-  4. (4)
-    $\operatorname{G-dim}_{R}(\omega)<\infty$.
-  5. (5)
-    $\operatorname{G-dim}_{R}(\omega^{*})<\infty$.
-    Now consider the following statements provided that $\operatorname{Ext}^{i}_{R}(\omega,R)=0$ for all $1\leq i\leq t$.
-  6. (6)
-    $\operatorname{Gid}_{R}((\omega^{*})^{\dagger})<\infty$.
-  7. (7)
-    $\operatorname{Gid}_{R}(\omega^{*})<\infty$.
-  8. (8)
-    $\operatorname{G-dim}_{R}((\omega^{*})^{\dagger})<\infty$.
-Then $\rm(1)-(5)$ are mutually equivalent, and each of them is equivalent to $\rm(6)$. In addition, $(6)\Rightarrow(7)\Leftrightarrow(8)$.
-Proof.
-The equivalences $(1)\Leftrightarrow(2)$ and $(3)\Leftrightarrow(4)$ follow directly from Corollary 4.2.
-Also, Theorem 5.6 asserts immediately that $(3)\Leftrightarrow(5)$.
-The equivalence $(1)\Leftrightarrow(3)$ is proved in [22, Theorem 2.1].
-Now, note that from the vanishing hypothesis and [21, Theorem 2.1(i)] we see that $\omega^{*}$ is maximal Cohen-Macaulay.
-Therefore $(5)\Leftrightarrow(6)$ and $(7)\Leftrightarrow(8)$ follow from Corollary 4.2.
-Since $(1)\Leftrightarrow(6)$, it is clear that $(6)\Rightarrow(7)$.
-∎
+Ghosh and Puthenpurakal proved in [15, Theorem 3.5] that if $M$ is a finitely generated $R$-module that is $\operatorname{G}$-perfect (i.e., $\operatorname{grade}(M)=\operatorname{G\text{-}dim}_{R}(M)$), then $\operatorname{r}(M)=\mu(\operatorname{Ext}^{t-r}_{R}(M,R))\operatorname{r}(R)$.
+A more general formula was later obtained in [32, Theorem (A)].
+As an application of Corollary 4.9, one can prove the Ghosh–Puthenpurakal formula to finitely generated $R$-modules $M$ over a Cohen–Macaulay ring RRsatisfying $\operatorname{G\text{-}dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $t-s\leq i\leq t-r$.
+Indeed, as in the proof of Corollary 4.10, this finiteness condition holds for all $i\geq 0$.
+Hence Corollary 4.9 yields $\mu(\operatorname{Ext}_{R}^{i}(M,\omega))=\mu(\operatorname{Ext}_{R}^{i}(M,R))\operatorname{r}(R)$, and the formula follows from [12, Theorem 3.1].
 5. Some questions
 In this section, we discuss some natural questions motivated by the results established earlier.
-Our primary question stems from the challenge of extending Corollaries 5.7 and 5.8 to the non-Cohen-Macaulay setting, removing the requirement that $R$ be a Cohen-Macaulay ring.
-Question 6.1.
+Our primary question stems from the challenge of extending Corollary 4.7 to the non-Cohen-Macaulay setting.
+Question 5.1.
 Let $M$ and $N$ be non-zero $R$-modules such that $\operatorname{pd}_{R}(N)<\infty$. Do the following facts hold?
@@
   2. (2)
-    If $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(M,N))<\infty$ and $\operatorname{Ext}_{R}^{i}(M,N)=0$ for all $1\leq i\leq t$, then $\operatorname{H-dim}_{R}(M)=0$.
-The following proposition partially answers Question 6.1 under the assumption that $\operatorname{H\text{-}dim}_{R}=\operatorname{pd}_{R}$.
+    If $\operatorname{H-dim}_{R}(\operatorname{Hom}_{R}(M,N))<\infty$ and $\operatorname{Ext}_{R}^{1\leq i\leq t}(M,N)=0$, then $\operatorname{H-dim}_{R}(M)=~0$.
+The following proposition partially answers Question 5.1 when $\operatorname{H\text{-}dim}_{R}=\operatorname{pd}_{R}$.
 Proposition 5.2.
@@
   2. (2)
-    $M$ has finite Gorenstein dimension and $s=t-0pt(M)$
+    $M$ has finite Gorenstein dimension equal to $s$.
 Proof.
-In both cases, the implication ()($\Rightarrow$)( ⇒ ) follows from Proposition 3.4. Thus, it remains to prove the implication ()($\Leftarrow$)( ⇐ ).
+In both cases, the implication $(\Rightarrow)$ follows from Corollary 2.3. Thus, it remains to prove the implication $(\Leftarrow)$.
 (1) Suppose that $\operatorname{pd}_{R}(N)<\infty$.
-By Theorem 2.1 and Lemma 5.4, $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(\Omega^{t}(M),N))<\infty$.
+By Lemma 4.4, we have $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(\Omega^{t}(M),N))<\infty$.
 Moreover, the vanishing condition shows that $\operatorname{Ext}_{R}^{i}(\Omega^{t}(M),R)=0$ for all $1\leq i\leq t$.
-Therefore, by [12, Proposition 19], $\operatorname{pd}_{R}(\Omega^{t}(M)^{\ast})<\infty$.
-Then by [12, Proposition 13], $\Omega^{t}(M)$ is free.
+Therefore, by [11, Proposition 19], $\operatorname{pd}_{R}(\Omega^{t}(M)^{\ast})<\infty$.
+Then by [11, Proposition 13], $\Omega^{t}(M)$ is free.
 Thus, $\operatorname{pd}_{R}(M)<\infty$.
@@
 ∎
-Remark 6.3.
-It should be noticed that in Proposition 6.2 the implication $\Rightarrow$ only requires that $\operatorname{pd}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for $0\leq i\leq t-\operatorname{depth}(M)$.
-For $N=R$, an answer to Question 6.1(2) is given by [12, Proposition 13].
-Next, we present an answer to Question 6.1(1) in the following proposition, which follows from and strengthens [12, Proposition 13] for projective dimension and Gorenstein dimension.
+For $N=R$, an answer to Question 5.1(2) is given by [11, Proposition 13].
+Next, we present an answer to Question 5.1(1) that strengthens this proposition.
 For that, we say that an $R$-module $M$ satisfies $(\widetilde{S}_{n})$ if $\operatorname{depth}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})\geq\min\{n,0pt_{R_{\mathfrak{p}}}(R_{\mathfrak{p}})\}$ for all prime ideals $\mathfrak{p}$ of $R$.
@@
 Proof.
-Let $s=t-n$.
-Then by Theorem 2.1 and Lemma 5.4, $\operatorname{H-dim}_{R}(\Omega^{s}(M)^{\ast})<\infty$.
+Let s=tns=t-n.
+Then by Lemma 4.4, $\operatorname{H-dim}_{R}(\Omega^{s}(M)^{\ast})<\infty$.
 Since $\operatorname{H-dim}_{R}(M^{\ast})<\infty$ and $M$ satisfies $(\widetilde{S}_{n})$, an application of the depth lemma shows that $\Omega^{s}(M)$ satisfies $(\widetilde{S}_{t})$.
-It follows from [12, Proposition 13] that $\operatorname{H-dim}_{R}(\Omega^{s}(M))=0$.
-Hence, by Proposition 2.3(3), $\operatorname{H-dim}_{R}(M)\leq s=t-n$.
+It follows from [11, Proposition 13] that $\operatorname{H-dim}_{R}(\Omega^{s}(M))=0$.
+Hence, by [4, Theorem 8.7(3)], $\operatorname{H-dim}_{R}(M)\leq s=t-n$.
 ∎
-Remark 6.5.
-It is important to mention Proposition 6.4 is significant only when $\operatorname{grade}(M)=0$ or $n=0$.
-The reason of this is that there are no non-zero $R$-modules $N$ with positive grade satisfying $(\widetilde{S}_{1})$.
-The proof of this fact is as follows.
-Suppose that $N$ satisfies $(\widetilde{S}_{1})$ and that $N^{\ast}=0$.
-Then, $\operatorname{pd}_{R}(\operatorname{Tr}_{R}(N))\leq 1$.
-As $N$ satisfies $(\widetilde{S}_{1})$, by [36, Theorem 5.8], we have that $\operatorname{Ext}_{R}^{1}(\operatorname{Tr}_{R}(N),R)=0$.
-Thus, $\operatorname{Tr}_{R}(N)$ is free and hence $N$ as well.
-But since $N^{\ast}=0,$ then $N=0$.
 Corollary 5.4.
 Let $M$ be an $R$-module. If $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(M,R))<\infty$ for all $0\leq i\leq t$, then $\operatorname{H-dim}_{R}(M)<\infty$.
-Note that Proposition 6.4 and Corollary 6.6 give some answers to [28, Question 4.11]. As an application of Corollary 6.6, we can improve [19, Proposition 3.3] as follows:
+Note that Proposition 5.3 and Corollary 5.4 give some answers to [24, Question 4.11]. As an application of Corollary 5.4, we can improve [16, Proposition 3.3].
 Corollary 5.5.
@@
 Then by Corollary 5.4, $\operatorname{H-dim}_{R}(M)<\infty$.
-Since $\operatorname{H-dim}_{R}(M)=\sup\{i\geq 0:\operatorname{Ext}_{R}^{i}(M,R)\not=0\}$ (see, for instance [26, Lemma 2.3(c)] together with [2, Theorem 1.4]) and $\operatorname{H-dim}_{R}(M)\leq t$, we see from the vanishing condition that $\operatorname{H-dim}_{R}(M)=h$.
+Since $\operatorname{H-dim}_{R}(M)=\sup\{i\geq 0:\operatorname{Ext}_{R}^{i}(M,R)\not=0\}$ (see, for instance [22, Lemma 2.3(c)] together with [2, Theorem 1.4]) and $\operatorname{H-dim}_{R}(M)\leq t$, we see from the vanishing condition that $\operatorname{H-dim}_{R}(M)=~h$.
 ∎
-The reader can note that Corollaries 5.7 and 5.8, which motivated Question 6.1, were obtained using Theorem 5.6, which was derived from Theorem 5.5.
-Thus, in view of the Foxby equivalence ([6, Theorem 3.3.2]), trying to answer these questions, we pose the following question:
-Question 6.8.
+The reader can note that Corollary 4.7, which motivated Question 5.1, was obtained using Theorem 4.6, which was derived from Theorem 4.5.
+Thus, in view of the Foxby equivalence ([6, Theorem 3.3.2]), we pose the following question:
+Question 5.6.
 It is possible to obtain a variation of Proposition 4.3 or Theorem 4.5 involving a dualizing complex $\boldsymbol{D}$ instead of $\omega$ or $C$?
 Now, motivated by Corollary 4.8, we pose the following:
-Question 6.9.
-Let $M$ and $N$ be non-zero $R$-modules such that $\operatorname{pd}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t$. Then does the equivalence
+Question 5.7.
+Let $M$ and $N$ be non-zero $R$-modules such that $\operatorname{pd}_{R}(\operatorname{Ext}_{R}^{i}(M,N))<\infty$ for all $0\leq i\leq t$ (or all $i\geq 0$). Then does the equivalence
 $\operatorname{id}_{R}(M)<\infty\Longleftrightarrow\operatorname{id}_{R}(N)<\infty$
 hold?
-One can note that the implication ()($\Leftarrow$)( ⇐ ) holds by Theorem 3.6.
+Note that the implication $(\Leftarrow)$ holds by Theorem 2.5.
 On the other hand, the other implication has a positive answer if $M$ is maximal Cohen-Macaulay.
@@
 Then $M\cong\omega^{n}$ for some $n>0$ and $\operatorname{pd}_{R}(\operatorname{Hom}_{R}(\omega,N))<\infty$, obtaining thus the desired conclusion from Proposition 4.1(3).
-Based on the blend between the Gorenstein dimensions, the canonical module, and its duals in Corollary 5.15, we close the paper by asking the following.
-Question 6.10.
-Assume $R$ is Cohen-Macaulay with canonical module $\omega$, and let $\_^{\dagger}$ stands for $\operatorname{Hom}_{R}(\_,\omega)$. Does any of the conditions
-  1. (1)
-    $\operatorname{Gid}_{R}(\omega^{*})<\infty$;
-  2. (2)
-    $\operatorname{G-dim}_{R}((\omega^{*})^{\dagger})<\infty$.
-imply that $R$ must be Gorenstein?
-Acknowledgment.
-This work was completed while the third author was visiting the University of Nebraska-Lincoln.
-He is grateful for the hospitality.
-He also thanks Roger and Sylvia Wiegand, Ryo Takahashi, Hailong Dao, Paulo Martins, Nawaj KC, Zachary Nason, Anna Brosowsky, Andrew Soto Levins and Kesavan Mohana Sundaram for some discussions about this work.
-In addition, he also thanks to Lars W.
-Christensen and Henrik Holm for their response to a question about completion and Gorenstein injective dimension.
-Additionally, the authors are grateful to Ryo Takahashi, Hailong Dao and Nawaj KC for suggesting some examples, and to Dipankar Ghosh for correcting some typos.
+Motivated by Corollary 3.3 and [24, Corollary 5.13], we close this section by asking the following.
+Question 5.8.
+Assume $R$ is Cohen-Macaulay with canonical module $\omega$, and let $^{\dagger}$ stand for $\operatorname{Hom}_{R}(\_,\omega)$.
+Does any of the conditions $\operatorname{Gid}_{R}(\omega^{*})<\infty$ or $\operatorname{G-dim}_{R}((\omega^{*})^{\dagger})<\infty$ imply that $R$ must be Gorenstein?
+Addendum
+After the first version of this paper, recently Questions 5.1 and 5.7 were affirmatively answered by Kaito Kimura in [20]. The technique in his paper differs from the one we used in Theorem 4.7.
+6. Examples
+We present some examples concerning Ext modules with finite homological dimension.
+Example 6.1.
+Assume $R$ is a local ring and let $\underline{x}=x_{1},\ldots,x_{n}$ be a regular sequence.
+Let $N$ be an $R$-module such that $\underline{x}N=0$.
+Set $M=R/(\underline{x})$.
+Since the minimal free resolution of $M$ is the Koszul complex of the sequence $\underline{x}$, we have
+$\operatorname{Ext}^{i}_{R}(M,N)\cong N^{\binom{n}{i}},\quad\forall\ i=0,\ldots,n.$
+Thus, $\operatorname{H-dim}_{R}(\operatorname{Ext}^{i}_{R}(M,N))<\infty$ for all $0\leq i\leq n$ if and only if $\operatorname{H-dim}_{R}(N)<\infty.$ Similarly, $\operatorname{Hid}_{R}(\operatorname{Ext}^{i}_{R}(M,N))<\infty$ for all $0\leq i\leq n$ if and only if $\operatorname{Hid}_{R}(N)<\infty.$
+Now, set $L=\oplus_{i=1}^{n}R/(x_{1},\ldots,x_{i})$ and let $X$ be a finitely generated $R$-module such that $\underline{x}$ is $X$-regular. For all $i=1,\ldots,n$, by the Koszul duality we must have
+$\operatorname{Ext}^{i}_{R}(L,X)\cong\operatorname{Ext}^{i}_{R}(R/(x_{1},\ldots,x_{i}),X)\cong X/(x_{1},\ldots,x_{i})X.$
+Thus $\operatorname{H-dim}_{R}(X)<\infty$ if and only if $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{i}(L,X))<\infty$ for all $0\leq i\leq n$.
+In particular, $\operatorname{H-dim}_{R}(\operatorname{Ext}^{i}_{R}(L,R))=i$, by the Auslander-Buchsbaum and Auslander-Bridger formulas.
+An $R$-module $M$ is said to be $\operatorname{H-}$perfect if $\operatorname{H-dim}_{R}(M)<\infty$ and $\operatorname{grade}(M)=\operatorname{H-dim}_{R}(M)$.
+Example 6.2.
+Let $M_{1},\ldots,M_{n}$ be H-perfect $R$-modules of grade $g_{1},\ldots,g_{n}$, respectively.
+Then each $\operatorname{Ext}_{R}^{g_{i}}(M_{i},R)$ have $\operatorname{H-dim}$ equal $g_{i}$ by [3, Theorem 6.3].
+Set $M=\bigoplus_{i=1}^{n}M_{i}$.
+Thus, for $j\not=g_{1},\ldots,g_{n}$, the $R$-module $\operatorname{Ext}_{R}^{j}(M,R)$ is zero, while that $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{g_{i}}(M,R))=g_{i}$ for each $1\leq i\leq n$.
+An $R$-module $M$ is $p$-spherical if $\operatorname{pd}_{R}(M)\leq p$ and $\operatorname{Ext}_{R}^{i}(M,R)=0$ for all $1\leq i\leq p-1$.
+Example 6.3.
+Based on the proof of [21, Proposition 6] we can construct spherical non-perfect modules as follows.
+We repeat the proof for convenience.
+Let $M$ be an $R$-module with positive grade $n$, and let $F_{\bullet}$ be a free resolution of $M$.
+Consider the truncation
+${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$${F_{n-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
+Set $N:=\operatorname{coker}(\phi^{*})$. Since $\operatorname{grade}(M)=n>0$, the complex
+${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{0}^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{1}^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{2}^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle$$\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${F_{n-1}^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi^{*}}$${F_{n}^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
+is a free resolution of $N$ and
+${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$${F_{n-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
+is exact.
+This means that $N$ is $n$-spherical as well as $N^{*}$ is a $(n+1)$th syzygy of $M$ (in the sense of [22]), and also $\operatorname{Ext}^{n}_{R}(N,R)\cong M$.
+It follows that whenever $M$ is such that $\operatorname{pd}_{R}(M)>\operatorname{grade}(M)=n>0$, there exists a $n$-spherical non-perfect $R$-module NN such that $\operatorname{Ext}^{n}_{R}(N,R)\cong M$.
+In particular, if $\operatorname{H-dim}_{R}(M)<\infty$ (resp.
+$\operatorname{Hid}_{R}(M)<\infty)$ then $\operatorname{H-dim}_{R}(\operatorname{Ext}_{R}^{n}(N,R))<\infty$ (resp.
+$\operatorname{Hid}_{R}(\operatorname{Ext}_{R}^{n}(N,R))<\infty$).
+Example 6.4.
+Let $k$ be a field and consider $R=k[[x,y]]/(xy)$. Then $M=R/(x)$ has infinite free resolution
+$\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${R\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{x}$${R\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{y}$${R\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{x}$${R\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
+so that $\operatorname{Ext}^{2i+1}_{R}(M,M)\cong M$ for all $i$, which means that $\operatorname{G-dim}_{R}(\operatorname{Ext}^{2i+1}_{R}(M,M))=0$ and $\operatorname{pd}_{R}(\operatorname{Ext}^{2i+1}_{R}(M,M))=\infty$.